### 09 Mar 2026 v1

# QED corrections of orders $m\alpha^{6}$ and $m\alpha^{6}(m/M)$ for HD+ rovibrational transitions beyond Born-Oppenheimer approximation

Abstract
The effective Hamiltonian of $m\alpha^{6}$ and $m\alpha^{6}(m/M)$ order corrections for hydrogen molecular ions has been derived in [ Z.-X. Zhong, et al., Phys. Rev. A 98, 032502(2018).], in this work we express the energy correction in the form of finite-value effective operators. The cut-off regularization scheme is used to determine finite part of divergent operators of the leading-order recoil corrections. Numerical calculations of first-order contributions are performed in the Hylleraas basis set. Combining the second-order terms calculated in recent work [V. I. Korobov, et al., Mol. Phys. e2563023 (2025).], the $m\alpha^{6}$-order corrections for the fundamental rovibrational transition are obtained with an uncertainty three times smaller than in previous calculations.
I introduction
Spectroscopy of hydrogen molecular ion HD+ has been used to determine proton-to-electron mass ratio [1, 2, 3, 4] and relative fundamental constants [5, 6], where HD+ was trapped in the Lamb-Dick region and its spectroscopy reached a precision of $10^{-11}$ or even higher. The corresponding theoretical calculations has been last for two decades, started from the non-relativistic energies [7, 8, 9] and leading-order relativistic and radiative corrections [10, 11, 12, 13, 14, 15], followed by $m\alpha^{6}$ order correction based on the Born-Oppenheimer approximation [16, 17] and $m\alpha^{7}$ order corrections [18, 19, 20]. In 2017, the theoretical spectroscopy reached a precision of $10^{-12}$ level by including adiabatic correction to the $m\alpha^{6}$ order contribution and the main contribution of $m\alpha^{8}$ order correction [21]. An update result adopting fundamental constants from CODATA 2018 [22] was presented in 2021 [23].
In the review of all the calculations, the non-relativistic energies and leading-order relativistic and ratiative corrections has been independent verified, while corrections of $m\alpha^{6}$ order and higher performed by only one group. In 2018, Zhong and coworkers derived spin-averaged effective Hamiltonian of orders $m\alpha^{6}$ and $m\alpha^{6}(m/M)$ for hydrogen molecular ions [24], which made an independent verification to HD+ $m\alpha^{6}$ order calculations [16, 17, 21] possible. It should be noted that the the recoil part contained some inaccuracies in derivation of relativisitic correction in Ref. [24],which was corrected recently in [25, 26]. In this work, we intend to calculate the first-order contribution of $m\alpha^{6}$ order relativistic corrections. After that, the complete spin-independent $m\alpha^{6}$ order contribution can be obtained after inclusion of the second-order contributions calulcated in [27].
This article is organized as follows: section II presents effective Hamiltonian, section III shows procedure of numerical calculations for the first-order and second-order perturbation contributions, numerical results and a conclusion are provided in section IV. Atomic units ($\hbar=m_{e}=e=1$) are used throughout this work, unless otherwise stated. Physical constants of CODATA 2018 [22] are adopted.
II the $m\alpha^{6}$ order effective Hamiltonian
In NRQED theory, an energy level of a light atomic system can be expanded in powers of the fine structure constant $\alpha$:
$E(\alpha)=E^{(2)}+E^{(4)}+E^{(5)}+E^{(6)}+E^{(7)}+O(\alpha^{8}),$ (1)
where $E^{(n)}$ represents a contribution of order $m\alpha^{n}$ and may include nuclear recoil terms. Each term $E^{(n)}$ can be obtained from the expectation value of the corresponding effective Hamiltonian.
In Eq. (1), $E^{(2)}\equiv E_{0}$ is the eigenvalue of the nonrelativistic Hamiltonian $H^{(2)}\equiv H_{0}$ with the associated eigenstate $\psi_{0}$ satisfying $H_{0}\psi_{0}=E_{0}\psi_{0}$, where
$H_{0}=\frac{\mathbf{p}_{e}^{2}}{2m_{e}}+\frac{\mathbf{p}_{1}^{2}}{2m_{1}}+\frac{\mathbf{p}_{2}^{2}}{2m_{2}}+\frac{z_{e}z_{1}}{r_{1}}+\frac{z_{e}z_{2}}{r_{2}}+\frac{z_{1}z_{2}}{r_{12}}.$ (2)
The Hamiltonian $H_{0}$ is expressed in the center-of-mass frame with the constraint $\mathbf{p}_{1}+\mathbf{p}_{2}+\mathbf{p}_{e}\equiv 0$. The relative position vectors between particles are defined as
$\mathbf{r}_{1}=\mathbf{r}_{e}-\mathbf{R}_{1},\quad\mathbf{r}_{2}=\mathbf{r}_{e}-\mathbf{R}_{2},\quad\mathbf{r}_{12}=\mathbf{R}_{2}-\mathbf{R}_{1},$ (3)
where $\mathbf{r}_{e}$, $\mathbf{R}_{1}$, and $\mathbf{R}_{2}$ denote the position vectors of the electron and nuclei 1 and 2, respectively. For convenient, the Coulomb potential is expressed in terms of pairwise interactions
$V=V_{1}+V_{2}+V_{12}\,$, (4)
where $V_{i}=z_{e}z_{i}/r_{i}$ and $V_{12}=z_{1}z_{2}/r_{12}$.
The term $E^{(4)}$ in Eq. (1) corresponds to the expectation value of the total Breit–Pauli Hamiltonian $H^{(4)}$. The second-order perturbation of $H^{(4)}$ contributes to the spin-averaged $m\alpha^{6}$-order correction. Since we are interested in the nonrecoil and leading-order recoil corrections of $H^{(6)}$, we consider only the required Hamiltonians:
$H^{(4)}$ $=H_{B}+H_{R}+H_{S},$ (5)
$H_{B}$ $=-\frac{\mathbf{p}_{e}^{4}}{8m_{e}^{3}}-\frac{\rho_{1}+\rho_{2}}{8m_{e}^{2}},$ (6)
$H_{R}$ $=\sum_{a}-\frac{z_{e}z_{a}}{m_{e}m_{a}}p_{e}^{i}\mathcal{W}^{ij}_{a}p_{a}^{j},$ (7)
$H_{S}$ $=\sum_{a}-z_{e}z_{a}\left[\frac{1+2a_{e}}{2m_{e}^{2}}\frac{1}{r_{a}^{3}}(\mathbf{r}_{a}\times\mathbf{p}_{e})-\frac{1+a_{e}}{m_{e}m_{a}}\frac{1}{r_{a}}(\mathbf{r}_{a}\times\mathbf{p}_{a})\right]\cdot\mathbf{s}_{e},$ (8)
where $a$ runs over the two nuclei, $\rho_{a}=-\Delta V_{a}=4\pi z_{e}z_{a}\delta(\vec{r}_{a})$, $\mathcal{W}^{ij}_{a}=(\delta^{ij}+r_{a}^{i}r_{a}^{j}/r_{a}^{2})/(2r_{a})$, and $a_{e}$ represents the electron’s anomalous magnetic moment.
The focus of this paper is on the $m\alpha^{6}$ order correction, denoted as $E^{(6)}$. This energy correction consists of several contributions of both order $m\alpha^{6}$ and order $m\alpha^{6}(m/M)$. The first set of contributions arises from the low-energy NRQED Hamiltonian [24] :
$H_{\text{rel}}^{(6)}$ $=\sum_{i=1}^{6}\delta H_{i},$ (9)
$\delta H_{1}$ $=\frac{\mathbf{p}_{e}^{6}}{16m_{e}^{5}},$ (10)
$\delta H_{2}$ $=\frac{1}{128m_{e}^{4}}[\mathbf{p}_{e}^{2},[\mathbf{p}_{e}^{2},V]]+\frac{3}{64m_{e}^{4}}\left\{\mathbf{p}_{e}^{2},\rho_{1}+\rho_{2}\right\},$ (11)
$\delta H_{3}$ $=\sum_{a}-\frac{1}{32m_{e}^{3}m_{a}}[\mathbf{p}_{a}^{2},[\mathbf{p}_{e}^{2},V_{a}]],$ (12)
$\delta H_{4}$ $=-\frac{1}{8m_{e}^{2}}\{\mathbf{p}_{e}^{2},2H_{R}\},$ (13)
$\delta H_{5}$ $=\sum_{a}\frac{1}{2m_{a}}\bigg\{\left[\frac{z_{e}z_{a}}{2r_{a}}\left(\delta^{ij}+\frac{r_{a}^{i}r_{a}^{j}}{r_{a}^{2}}\right)\frac{p_{e}^{j}}{m_{e}}\right]$
$\times\left[\frac{z_{e}z_{a}}{2r_{a}}\left(\delta^{il}+\frac{r_{a}^{i}r_{a}^{l}}{r_{a}^{2}}\right)\frac{p_{e}^{l}}{m_{e}}\right]+\frac{1}{2m_{e}^{2}}\bm{\varepsilon}_{a}^{2}\bigg\},$ (14)
$\delta H_{6}$ $=\sum_{a}-\frac{1}{m_{e}m_{a}}\left\{[p_{e}^{i},V]\mathcal{X}^{ij}_{a}[V,p_{a}^{j}]+p_{e}^{i}[\mathcal{X}^{ij}_{a},\frac{\mathbf{p}_{e}^{2}}{2m_{e}}][V,p_{a}^{j}]\right\},$ (15)
where $\bm{\varepsilon}_{a}=-\bm{\nabla}V_{a}$ and $\mathcal{X}^{ij}_{a}=\left[r^{i}_{a}r^{j}_{a}-3\delta^{ij}r^{2}_{a}\right]/(8r_{a})$. Curly brackets denote $\{A,B\}=AB+B^{*}A^{*}$. Another contribution — the pure recoil correction — is determined by the high-energy region as the contact type interaction of the NRQED [28]:
$H_{H}^{(6)}=\sum_{a}\left(4\ln 2-\frac{7}{2}\right)\frac{z_{a}^{3}}{m_{a}}\pi\delta(\mathbf{r}_{a}).$ (16)
Formally, the contribution of the three-photon exchange diagrams in (16) diverges, but, as it was shown [29], it is cancelled out by the corresponding term from the low-energy region.
The second set of contributions to $E^{(6)}$ arises from radiative corrections of orders $m\alpha^{6}$ and $m\alpha^{6}(m/M)$, denoted by $H_{\text{rad}}^{(6)}$ and $H_{\text{rad-rec}}^{(6)}$, respectively. These corrections are derived within the bound-state QED. The $m\alpha^{6}$-order nonrecoil contribution includes both one-loop and two-loop corrections and may be written in the form: [30, 31]
$\begin{array}[]{@{}l}\displaystyle H_{\text{rad}}^{(6)}=\frac{4\pi}{m_{e}^{2}}\bigg\{\left(\frac{139}{128}-\frac{\ln 2}{2}+\frac{5}{192}\right)z_{a}^{2}\delta(\mathbf{r}_{a})\\[8.53581pt] \displaystyle\hskip 28.45274pt-\frac{1}{4\pi^{2}}\left[\frac{2179}{648}+\frac{10}{27}\pi^{2}-\frac{3}{2}\pi^{2}\ln 2+\frac{9}{4}\zeta(3)\right]z_{a}\delta(\mathbf{r}_{a})\bigg\}.\end{array}$ (19)
The radiative recoil high-energy contribution of the order of $m\alpha^{6}(m/M)$ is determined by the following expression [31, 32]:
$H_{\text{rad-rec}}^{(6)}$ $=\frac{m_{e}}{m_{a}}\bigg\{\left[\frac{3}{4}+\frac{6}{\pi^{2}}\zeta(3)-\frac{14}{\pi^{2}}-2\ln 2\right]z_{a}^{2}\pi\delta(\mathbf{r}_{a})$
$+\left(\frac{2}{9}\pi^{2}-\frac{70}{27}\right)\frac{z_{a}^{2}}{\pi}\delta(\mathbf{r}_{a})\bigg\}.$ (20)
And the final set of contributions to $E^{(6)}$ consists of second-order perturbation corrections $E_{\text{sec}}$ due to the Breit–Pauli Hamiltonian $H^{(4)}$:
$E_{\text{sec}}^{(6)}$ $=E_{B}+E_{R}+E_{S},$ (21)
[…]
where $E_{B}$ is the contribution from the Breit interaction, $E_{R}$ is the relativistic correction due to retardation effects, $E_{S}$ is the contribution from spin–orbit interactions and $|Q=1-\outerproduct{\psi_{0}}{\psi_{0}}$ is the projection operator that excludes the reference state $\psi_{0}$.
The total $m\alpha^{6}$-order correction is obtained by summing the contributions from Eqs. (9), (16), (19), (20), and (21):
$E^{(6)}=\expectationvalue{H_{\text{rel}}^{(6)}+H_{H}^{(6)}+H_{\text{rad}}^{(6)}+H_{\text{rad-rec}}^{(6)}}+E_{\text{sec}}^{(6)}.$ (25)
Both the first-order perturbation contribution (the terms within angle brackets in Eq. (25)) and the second-order contribution $E_{\text{sec}}^{(6)}$ exhibit native singularities due to the perturbation theory framework, which can and should be regularized. The details of this regularization are discussed in the next section.
III Elimination of Singularities and Regularization
In quantum electrodynamics (QED), various regularization schemes are employed to handle singularities in theoretical calculations. As recently demonstrated by Korobov [25, 26], all these regularization methods yield unified finite expressions that are computationally tractable.
In our approach, we adopt the coordinate-space cutoff regularization scheme, where divergent integrals are rendered finite by introducing a minimal radial cutoff parameter $r_{0}$ as the lower integration limit. The complete mathematical formulation and implementation details of this coordinate cutoff regularization procedure are thoroughly discussed in Refs. [25, 26].
The effective Hamiltonian that requires regularization is the relativistic correction $H_{\text{rel}}$ of order $m\alpha^{6}$ in Eq. (9). Additionally, the second-order corrections $E_{B}$ and $E_{R}$ given by Eqs. (22) and (23), respectively, must be addressed.
As can be seen from Appendix C, singularities of the effective Hamiltonian can be isolated in the form of singular integrals of three types: $\langle V_{a}\rho_{a}\rangle$, $\langle V_{a}^{3}\rangle$ and $\langle\bm{\varepsilon}_{a}^{2}\rangle$. Then $\langle V_{a}\rho_{a}\rangle$ is transferred into the other two singular integrals using the identity (86) while $\langle V_{a}^{3}\rangle$ and $\langle\bm{\varepsilon}_{a}^{2}\rangle$ can be redefined as finite distributions using the coordinate cutoff regularization, see Eq. (70). In the following text, superscript $\langle\cdot\rangle^{(m)}$ denotes that the operator in the brackets is modified according to these prescriptions. In the nonrecoil correction, these singularities disappear from the effective Hamiltonians, since they cancel each other out. Thus, superscript $\langle\cdot\rangle^{\prime}$ is used to indicate that regularized singular integrals of the above type are removed from the matrix element $\langle\cdot\rangle^{(m)}$.
In this section we consider only the two leading orders of the expansion in $m_{e}/m_{a}$, the higher order terms are ignored.
III.1 Singularity in nonrecoil correction
It has been demonstrated that the singularities in the $m\alpha^{6}$-order nonrecoil correction cancel out exactly in the hydrogen case [16, 25]. For hydrogen molecular ions, the first-order perturbation contributions correspond to the effective Hamiltonians $\delta H_{1}$$\delta H_{3}$ (see Eqs. (10)–(12)). The singular part of the first-order contribution can be isolated as (see Eqs.(66)-(68) of Ref. [24]):
$S_{1,2,3}=\sum_{a}\bigg[\frac{1}{32m_{e}^{3}}\left(1-\frac{5m_{e}}{m_{a}}\right)\langle\bm{\varepsilon}_{a}^{2}\rangle-\frac{1}{8m_{e}^{2}}\left(1-\frac{6m_{e}}{m_{a}}\right)\langle V_{a}^{3}\rangle\bigg]\,.$ (26)
The nonrecoil second-order contribution is given by $E_{B}$ (22) and $E_{S}$ (24). To isolate the singularity in $E_{B}$, we introduce the transformation:
$H_{B}=H^{\prime}_{B}+(E_{0}-H_{0})U+U(E_{0}-H_{0})\,$, (27)
where
$U=\lambda_{1}V_{1}+\lambda_{2}V_{2}\,,\quad\lambda_{a}=-\frac{1}{4m_{e}}\left(1-\frac{3m_{e}}{m_{a}}\right)\,.$ (28)
This allows $E_{B}$ to be expanded as:
[…]
$-2\langle U\rangle\langle H_{B}\rangle-\langle U(E_{0}-H_{0})U\rangle\,$, (29)
where the first term ($E^{\prime}_{B}$) is numerically tractable, while the remaining three terms constitute a first-order contribution
$H_{B}^{\text{1st}}=\{H_{B},U\}-2\langle{H_{B}}\rangle U-U(E_{0}-H_{0})U\,.$ (30)
The singular part of $H_{B}^{\text{1st}}$ can be extracted as $S_{B}$ :
$S_{B}=-\sum_{a}\bigg[\frac{1}{32m_{e}^{3}}\left(1-\frac{5m_{e}}{m_{a}}\right)\langle\bm{\varepsilon}_{a}^{2}\rangle-\frac{1}{8m_{e}^{2}}\left(1-\frac{6m_{e}}{m_{a}}\right)\langle V_{a}^{3}\rangle\bigg]\,.$ (31)
One can see that singularities in the nonrecoil correction after summation of all contributions are cancelled out $S_{1,2,3}+S_{B}\equiv 0$.
Once singularities are removed, the modified effective Hamiltonian of the nonrecoil relativistic corrections is as
${E^{\prime}}_{\text{rel}}^{(6)}=\langle\delta H^{\prime}_{1}+\delta H^{\prime}_{2}+\delta H^{\prime}_{3}+{H^{\prime}}_{B}^{\text{1st}}\rangle\,.$ (32)
The subtraction of the singularities and the determination of the finite expectation values for modified operators are moved to the Appendix; see sections B and C for details. Here we present the final results only:
[…]
and
$\begin{array}[]{@{}l}\displaystyle\langle{H^{\prime}}_{B}^{\text{1st}}\rangle=-2\langle H_{B}\rangle\langle U\rangle+\frac{\lambda_{1}\lambda_{2}}{m_{e}}\langle\bm{\varepsilon}_{1}\bm{\varepsilon}_{2}\rangle\\[5.69054pt] \displaystyle-\sum_{a}\frac{\lambda_{a}}{8m_{e}^{3}}\left(\langle\{p_{e}^{4},V_{a}\}\rangle^{\prime}\!+\!2m_{e}\langle V_{a}\rho_{b}\rangle\!+\!2m_{e}\langle V_{a}\rho_{a}\rangle^{\prime}\right)\,.\end{array}$ (40)
For convenience, we may not expand all the matrix elements into explicit formulas.
III.2 Singularity in recoil correction
We now proceed to isolate the singularities in the $m\alpha^{6}(m/M)$-order recoil correction, Eqs. (13)-(15) and (16). It should be noted that our previous consideration of this part [24] was incomplete. Here we will use the results of Ref. [26], where it was shown that the singularities of recoil corrections from the low-energy region contribution are cancelled out when the singularity of the high-energy contribution $H_{H}^{(6)}$ (16) is taken into account. The correct results are presented below; the details can be found in Appendix C.
Let us first consider $\delta H_{4}$, which can be explicitly expressed as
$\delta H_{4}$ $=-\frac{1}{2m_{e}}\{E_{0}-V,H_{R}\}$
$=-\frac{E_{0}H_{R}}{m_{e}}-\sum_{a}\frac{z_{e}z_{a}}{2m_{e}^{2}m_{a}}\{V,p_{e}^{i}\mathcal{W}^{ij}_{a}p_{a}^{j}\}\,$, (41)
where the main part of the second term can be expanded as
$\langle\{V,p_{e}^{i}\mathcal{W}^{ij}_{a}p_{a}^{j}\}\rangle=2\langle p_{e}^{i}\mathcal{W}^{ij}_{a}Vp_{a}^{j}\rangle-\langle[p_{e}^{i},V]\mathcal{W}^{ij}_{a}p_{a}^{j}\rangle$
$\qquad+\langle p_{e}^{i}\mathcal{W}^{ij}_{a}[p_{a}^{j},V]\rangle$
$=2\langle p_{e}^{i}\mathcal{W}^{ij}_{a}Vp_{a}^{j}\rangle-\langle\imath\varepsilon_{a}^{i}\mathcal{W}_{a}^{ij}p_{a}^{j}\rangle-\langle\imath\varepsilon_{b}^{i}\mathcal{W}_{a}^{ij}p_{a}^{j}\rangle_{b\neq a}$
$\qquad-\langle p_{e}^{i}\mathcal{W}_{a}^{ij}\imath\varepsilon_{a}^{j}\rangle\mp\langle p_{e}^{i}\mathcal{W}_{a}^{ij}\imath\bm{\varepsilon}_{12}^{j}\rangle$
$=2\langle p_{e}^{i}\mathcal{W}^{ij}_{a}Vp_{a}^{j}\rangle-\langle\imath\varepsilon_{a}^{i}\mathcal{W}_{a}^{ij}p_{a}^{j}\rangle-\langle p_{e}^{i}\mathcal{W}_{a}^{ij}\imath\varepsilon_{a}^{j}\rangle\,.$ (42)
Numerical calculations confirm that the matrix elements $\langle\imath\varepsilon_{b}^{i}\mathcal{W}_{a}^{ij}p_{a}^{j}\rangle_{b\neq a}$ and $\langle p_{e}^{i}\mathcal{W}_{a}^{ij}\imath\bm{\varepsilon}_{12}^{j}\rangle$ are small and may be neglected. The first term in the last line of Eq. (III.2) is regular, while the other two are singular. And the singularity in $\delta H_{4}$ can be written as
$S_{4}=-\sum_{a}\frac{1}{2m_{e}^{2}m_{a}}\left(\langle V_{a}\rho_{a}\rangle+\langle\bm{\varepsilon}_{a}^{2}\rangle\right)\,.$ (43)
For $\delta H_{5}$ one can directly obtain that the singular part may be written as
$S_{5}=\sum_{a}\frac{1}{4m_{e}^{2}m_{a}}\langle\bm{\varepsilon}_{a}^{2}\rangle\,.$ (44)
The derivation of the singular part in $\delta H_{6}$ is more complicated and we have moved it to the Appendix C. Here we present the final result only,
$S_{6}=\sum_{a}\frac{1}{8m_{e}^{2}m_{a}}\left(\langle\bm{\varepsilon}_{a}^{2}\rangle-2m_{e}\langle V_{a}^{3}\rangle+\langle V_{a}\rho_{a}\rangle\right)\,.$ (45)
Next we should consider the second-order contribution to the recoil correction: $E_{R}$ (23). Applying the transformation (27), $E_{R}$ takes the form
$E_{R}$ $=2\langle H^{\prime}_{B}Q(E_{0}-H_{0})^{-1}QH_{R}\rangle+\langle\{H_{R},U\}\rangle-2\langle H_{R}\rangle\langle U\rangle\,$, (46)
where the first term ($E^{\prime}_{R}$) is regular and the other two terms can be thought of as a modification of the effective Hamiltonian and considered as the first-order contribution
$H_{R}^{\text{1st}}=\{H_{R},U\}-2\langle H_{R}\rangle U\,.$ (47)
The first term $H_{R}^{\text{1st}}$ is singular and can be considered in a similar way as $\delta H_{4}$. Thus, the singular part of $E_{R}$ is
$S_{R}=\sum_{a}\frac{1}{4m_{e}^{2}m_{a}}\left(\langle\bm{\varepsilon}_{a}^{2}\rangle+\langle V_{a}\rho_{a}\rangle\right)\,.$ (48)
Collecting all the singularities in Eqs. (43), (44), (45) and (48), one gets
$S_{\text{rec}}$ $=S_{4}+S_{5}+S_{6}+S_{R}$
$=\sum_{a}\frac{1}{m_{e}^{2}m_{a}}\left(\frac{\langle\bm{\varepsilon}_{a}^{2}\rangle}{8}-\frac{m_{e}\langle V_{a}^{3}\rangle}{4}-\frac{\langle V_{a}\rho_{a}\rangle}{8}\right)\,.$ (49)
Therefore, we should study integrals of those matrix elements $\langle\bm{\varepsilon}_{a}^{2}\rangle$, $\langle V_{a}^{3}\rangle$ and $\langle V_{a}\rho_{a}\rangle$, whose singular parts have been explicitly separated in Ref. [25] and Appendix A.
$S_{\text{rec}}$ will be regularized as $R_{\text{rec}}$, namely regular part of $S_{\text{sec}}$, by eliminating singularity of high-energy contribution $H_{H}^{(6)}$. Subsequently, the regularization of recoil contribution will rise a term determined by the $\delta$-function distribution. In the cut-off regularization, it is expressed as
$H_{\text{reg}}=\sum_{a}C_{r}\frac{z_{a}^{3}}{m_{a}}\pi\delta(\mathbf{r}_{a})\,.$ (50)
The coefficient $C_{r}=4\ln 2-2$ is determined by comparing expectation value of regularized effective Hamiltonian with the anayltical total $m\alpha^{6}$-oder correction in hydrogen atom, see Appendix D.
After regularization of all the contributions at $m\alpha^{6}(m/M)$ order, we obtain the effective Hamiltonians for the recoil relativistic corrections as follows:
${E^{\prime}}^{(6)}_{\text{rel-rec}}=\langle\delta H^{\prime}_{4}+\delta H^{\prime}_{5}+\delta H^{\prime}_{6}+R_{\text{rec}}+H_{\text{reg}}+{H^{\prime}}^{\text{1st}}_{R}\rangle\,$, (51)
where
$\langle\delta H^{\prime}_{4}\rangle$ $=-\frac{E_{0}}{m_{e}}\langle H_{R}\rangle-\sum_{a}\frac{z_{e}z_{a}}{2m_{e}^{2}m_{a}}\langle 2p_{e}^{i}\mathcal{W}^{ij}_{a}Vp_{a}^{j}\rangle\,$, (52)
$\langle\delta H^{\prime}_{5}\rangle$ $=\sum_{a}\bigg[\frac{1}{8m_{e}^{2}m_{a}}\bigg(\langle\mathbf{p}_{e}V_{a}^{2}\mathbf{p}_{e}\rangle+3\left\langle(\mathbf{p}_{e}\cdot\mathbf{r}_{a})\frac{V_{a}^{2}}{r_{a}^{2}}(\mathbf{r}_{a}\cdot\mathbf{p}_{e})\right\rangle\bigg)\,$, (53)
$\langle\delta H^{\prime}_{6}\rangle$ $=\sum_{a}\frac{1}{m_{e}m_{a}}\bigg\{\frac{1}{8m_{e}}\bigg[\left\langle\mathbf{p}_{e}V_{a}^{2}\mathbf{p}_{e}\right\rangle$
$\quad-3\left\langle(\mathbf{p}_{e}\cdot\mathbf{r}_{a})\frac{V_{a}^{2}}{r_{a}^{2}}(\mathbf{r}_{a}\cdot\mathbf{p}_{e})\right\rangle\bigg]-\frac{z_{e}z_{a}}{4}\langle r_{a}\bm{\varepsilon}_{a}\bm{\varepsilon}_{b}\rangle$
$\quad\pm z_{e}z_{a}\bigg\langle[p_{e}^{i},V]\mathcal{X}^{ij}_{a}\imath\bm{\varepsilon}_{12}^{j}\bigg\rangle\bigg\}\,$, (54)
and
$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle$ $=\sum_{a,b\neq a}\frac{z_{e}z_{a}}{2m_{e}^{2}m_{a}}\langle p_{e}^{i}\mathcal{W}_{a}^{ij}Vp_{a}^{j}\rangle-2\langle H_{R}\rangle\langle U\rangle\,.$ (55)
It should be noted that the regular part $R_{\text{rec}}$ of the divergent matrix elements appeared in $S_{\text{rec}}$ (III.2) should be retained after the regularization.
III.3 the regularized effective Hamiltonian
We now sum all contributions from Eqs. (16), (19), (20), (24), (32), (50) and (51) to obtain the total corrections at orders $m\alpha^{6}$ and $m\alpha^{6}(m/M)$:
$E^{(6)}$ $={E^{\prime}}_{\text{rel}}^{(6)}+{E^{\prime}}_{\text{sec}}^{(6)}+E_{\text{rad}}^{(6)}+{E^{\prime}}_{\text{rel-rec}}^{(6)}+E_{\text{rad-rec}}^{(6)}\,.$ (56)
Comparing with previous $E^{(6)}$ of Eq. (25), the relativistic correction due the effective Hamiltonian $H_{\text{rel}}^{(6)}$ has been regularized and split into nonrecoil correction $E_{\text{rel}}^{(6)}$ (32) and recoil correction $E_{\text{rel-rec}}^{(6)}$ (51). It should be noted that the high-energy relativistic correction, so called pure recoil correction $H_{H}^{(6)}$, has been embraced in the contribution arising in the regularization $H_{\text{reg}}$ (50). On the other hand, the second-order corrections ${E}_{\text{sec}}^{(6)}$ is regularized as ${E^{\prime}}_{\text{sec}}^{(6)}=E^{\prime}_{B}+E_{S}+E^{\prime}_{R}$ where $E^{\prime}_{B}=\langle H^{\prime}_{B}Q(E_{0}-H_{0})^{-1}QH^{\prime}_{B}\rangle$, $E^{\prime}_{R}=2\langle H^{\prime}_{B}Q(E_{0}-H_{0})^{-1}QH_{R}\rangle$ and $E_{S}$ (24) are regular enough for numerical study.
IV Numerical Calculations and Conclusion
Table 1: Values of regularied divergent matrix elements for HD+ rovibrational states (in atomic units). The numerical values given are exact to the last digit unless otherwise noted. Operators $\langle V_{1}\rho_{1}\rangle^{\prime}$ are defined as in (87).
(0,0) (0,1) (1,0) (1,1)
$\langle V_{1}^{3}\rangle_{r}$ 1.6947 2934 9762(2) 1.6601 8538 3679(7) 1.6932 1792 262(3) 1.6587 5603 832(2)
$\langle V_{2}^{3}\rangle_{r}$ 1.6913 2284 4951 1.6567 3745 644(2) 1.6898 1215 596(7) 1.6553 0853 833(3)
$\langle\varepsilon_{1}^{2}\rangle_{r}$ -1.4268 3109 35(2) –1.4012 5747 88(6) –1.4256 45901(2) –1.4001 32322(2)
$\langle\varepsilon_{2}^{2}\rangle_{r}$ -1.4254 5127 0134(7) –1.3998 02588(1) –1.4242 65356(5) –1.3986 76515(2)
$\langle V_{1}\rho_{1}\rangle^{\prime}$ –2.9993 4511 0567 –2.9150 8028 6650 –2.996 5836 6165 –2.9124 8207 422(6)
$\langle V_{2}\rho_{2}\rangle^{\prime}$ –2.9948 1780 0249 –2.9104 8557 9425 –2.992 0570 7310 –2.9078 8765 6271(8)
$\langle V_{1}^{2}\mathbf{p}_{e}^{2}\rangle^{\prime}$ –1.4813 9479 1135 –1.4385 6298 1389 –1.4804 2441 1796 –1.4376 6199 5594
$\langle V_{2}^{2}\mathbf{p}_{e}^{2}\rangle^{\prime}$ –1.4788 9959 4264 –1.4360 3955 2156 –1.4779 2998 6340 –1.4351 3910 9269(5)
$\langle V_{1}^{2}\mathbf{p}_{1}^{2}\rangle^{\prime}$ 5.9701 0577 1673 20.3159 9546 6009 6.6687 9581 7755(3) 20.9624 8160 67(2)
$\langle V_{2}^{2}\mathbf{p}_{2}^{2}\rangle^{\prime}$ 5.9623 8727 4107 20.2891 4197 9807 6.6609 00330(6) 20.9354 8110(1)
$\langle V\mathbf{p}_{1}^{2}V\rangle^{\prime}$ 12.7118 9202 2087 39.2947 8595 0492 14.0043 4653 5632(3) 40.4898 3443 34(2)
$\langle V\mathbf{p}_{2}^{2}V\rangle^{\prime}$ 12.7102 4302 4887 39.2859 6578 8716 14.0030 73800(6) 40.4814 0269(1)
$\langle\{\mathbf{p}_{e}^{2},\rho_{1}\}\rangle^{\prime}$ 18.2471 9863 2286 17.7214 5445 3675 18.2305 9594 1969(2) 17.7058 8161 45(4)
$\langle\{\mathbf{p}_{e}^{2},\rho_{2}\}\rangle^{\prime}$ 18.2147 1435 2833 17.6887 0915 2486 18.1981 2087 512(5) 17.6731 4265 18(2)
$\langle\{\mathbf{p}_{e}^{4},V_{1}\}\rangle^{\prime}$ 8.0683 0063 2230 7.8427 3701 1450 8.0647 8568 0309(5) 7.8396 3767 03(2)
$\langle\{\mathbf{p}_{e}^{4},V_{2}\}\rangle^{\prime}$ 8.0437 1320 4651 7.8181 1053 5982 8.0402 1158 890(2) 7.8150 2238 49(1)
Table 2: Values of regular matrix elements for HD+ rovibrational states (in atomic units). The numerical values given are exact to the last digit unless otherwise noted.
(0,0) (0,1) (1,0) (1,1)
$E_{0}$ –0.5978 9796 8610 –0.5891 8182 9561 –0.5976 9812 8194 –0.5889 9111 1996
$\langle U\rangle$ 0.4214 4656 2271 0.4129 4440 6667 0.4212 2454 0596 0.4127 3112 3370
$\langle H_{B}\rangle$ –0.1366 0153 9235 –0.1338 7800 1186 –0.1364 6376 9526 –0.1337 4881 366(2)
$\langle H_{R}\rangle$ –0.0004 7880 1585 –0.0004 6806 9095 –0.0004 7839 1583 –0.0004 6768 4165
$\langle V_{1}\rho_{2}\rangle$ 1.2932 5736 5489 1.2445 4449 4788 1.2907 9418 6712(2) 1.2421 6853 875(1)
$\langle V_{2}\rho_{1}\rangle$ 1.2949 7113 2033 1.2462 3915 8675 1.2925 0636 9768 1.2438 6174 637(5)
$\langle\mathbf{p}_{1}^{2}\rangle$ 5.9805 4596 1465 16.0975 3333 1318 6.4633 5834 8141 16.5472 3819 2988
$\langle\mathbf{p}_{2}^{2}\rangle$ 5.9782 9548 7288 16.0925 0308 7347 6.4609 7631 9844 16.5420 8531 8687
$\langle V\mathbf{p}_{1}^{2}\rangle$ –8.3947 4694 8040 –20.3714 7544 2789 –8.9728 6271 9712 –20.9059 8595 098(5)
$\langle V\mathbf{p}_{2}^{2}\rangle$ –8.3901 0978 4358 –20.3635 0767 2653 –8.9680 6789 950(3) –20.8978 7211 8122(5)
$\langle\mathbf{p}_{e}^{4}\rangle$ 6.3001 9994 7706 6.1590 2235 2359 6.2944 5074 6088(5) 6.1535 9974 587(8)
$\langle\{\mathbf{p}_{e}^{4},V_{12}\}\rangle$ 6.2640 6896 7646 6.0337 6087 260(1) 6.2520 0879 745(2) 6.0221 1748 900(3)
$\langle\bm{\varepsilon}_{1}\cdot\bm{\varepsilon}_{2}\rangle$ –0.0544 9531 3576 –0.0508 0894 3618 –0.0543 7418 1182 –0.0506 9770 3844
$\langle\bm{\varepsilon}_{1}\cdot\bm{\varepsilon}_{12}\rangle$ –0.0615 4079 8892 –0.0590 6505 7552 –0.0613 4476 1159 –0.0588 7417 7983
$\langle\bm{\varepsilon}_{2}\cdot\bm{\varepsilon}_{12}\rangle$ 0.0615 4700 6305 0.0590 7129 4583 0.0613 5094 1177 0.0588 8038 6718
$\langle\mathbf{p}_{e}V_{1}^{2}\mathbf{p}_{e}\rangle$ 1.5179 5031 9432 1.4765 1730 5261 1.5161 5924 9854 1.4748 2007 862(6)
$\langle\mathbf{p}_{e}V_{2}^{2}\mathbf{p}_{e}\rangle$ 1.5159 1820 5985 1.4744 4602 7269 1.5141 2708 6759(5) 1.4727 4854 7002(6)
$\langle(\mathbf{p}_{e}\cdot\mathbf{r}_{1})(V_{1}^{2}/r_{1}^{2})(\mathbf{r}_{1}\cdot\mathbf{p}_{e})\rangle$ 1.3200 5514 4383 1.2895 5695 8897 1.3187 0656 3776 1.2882 8015 190(8)
$\langle(\mathbf{p}_{e}\cdot\mathbf{r}_{2})(V_{2}^{2}/r_{2}^{2})(\mathbf{r}_{2}\cdot\mathbf{p}_{e})\rangle$ 1.3178 7214 0481 1.2873 3731 9221 1.3165 2379 6039(5) 1.2860 6053 815(1)
$\langle z_{e}z_{1}p_{e}^{i}\mathcal{W}_{1}^{ij}Vp_{1}^{j}\rangle$ –1.2281 0539 4297 –1.2008 2524 9325 –1.2270 8820 5793 –1.1998 6987 832(5)
$\langle z_{e}z_{2}p_{e}^{i}\mathcal{W}_{2}^{ij}Vp_{2}^{j}\rangle$ –1.2241 3289 5621 –1.1936 4318 5135 –1.2229 7915 8165(5) –1.1925 6219 346(2)
$\langle z_{e}z_{1}r_{1}\bm{\varepsilon}_{1}\cdot\bm{\varepsilon}_{2}\rangle$ 0.0169 4925 9569 0.0148 8857 1832 0.0169 8530 2788 0.01492 6808 1686
$\langle z_{e}z_{1}r_{2}\bm{\varepsilon}_{1}\cdot\bm{\varepsilon}_{2}\rangle$ 0.0169 1867 6737 0.0148 5875 0542 0.0169 5477 8808 0.01489 7043 4046
$\langle z_{e}z_{1}[p_{e}^{i},V]\mathcal{X}_{1}^{ij}\imath\varepsilon_{12}^{j}\rangle$ 0.0516 8603 2319 0.0495 1598 2047 0.0515 5049 9805 0.0493 8474 3758
$\langle z_{e}z_{2}[p_{e}^{i},V]\mathcal{X}_{2}^{ij}\imath\varepsilon_{12}^{j}\rangle$ –0.0517 2674 9355 –0.0495 5592 2695 –0.0515 9115 4163 –0.0494 2462 4074
Table 3: Expectation values of modified effective Hamiltonians for HD+ rovibrational states (in atomic units). The numerical values given are exact to the last digit unless otherwise noted.
(0,0) (0,1) (1,0) (1,1)
$\langle\delta H^{\prime}_{1}\rangle$ –1.8703 480(6) –1.8137 755(5) –1.8687 273(6) –1.8122 633(5)
$\langle\delta H^{\prime}_{2}\rangle$ 1.7125 566(5) 1.6630 255(5) 1.7109 929(5) 1.6615 589(5)
$\langle\delta H^{\prime}_{3}\rangle$ –0.0000 0628 5744(2) –0.0000 0603 2892(2) –0.0000 0626 5720(2) –0.0000 0601 3395(2)
$\langle\delta{H^{\prime}}_{B}^{\text{1st}}\rangle$ 0.4020 166(1) 0.3880 342(1) 0.4016 656(1) 0.3877 119(1)
${E^{\prime}}_{\text{rel}}^{(6)}$ 0.2442 189(8) 0.2372 782(7) 0.2439 249(8) 0.2370 001(7)
$\langle\delta H^{\prime}_{4}\rangle$ 0.0007 150(4) 0.0007 015(4) 0.0007 144(4) 0.0007 010(4)
$\langle\delta H^{\prime}_{5}\rangle$ 0.0005 5890 9980 0.0005 4532 3594 0.0005 5831 3896 0.0005 4475 9030
$\langle\delta H^{\prime}_{6}\rangle$ –0.0002 1032 7438 –0.0002 0656 2370 –0.0002 1021 536(2) –0.0002 0645 96(2)
$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle$ –0.0002 1544 0395 –0.0002 1617 5234(2) –0.0002 1533 761(1) –0.0002 1606 7579(5)
$\langle R_{\text{rec}}\rangle$ 0.0000 3966 5983 0.0000 3761 7674 0.0000 4008 5121 0.0000 3757 6478
$\langle H_{\text{reg}}\rangle$ 0.0004 1079 5750 0.0004 0137 2924 0.0004 1042 9068 0.0004 0102 6584
${E^{\prime}}_{\text{rel-rec}}^{(6)}$ 0.0012 986(4) 0.0012 630(4) 0.0017 977(4) 0.0012 618(4)
Table 4: QED corrections of order $m\alpha^{6}$ for HD+ rovibrational transitions $(L,v)\rightarrow(L^{\prime},v^{\prime})$, in kHz.
$(0,0)\rightarrow(0,1)$ $(0,0)\rightarrow(1,0)$ $(0,0)\rightarrow(1,1)$
$\Delta E_{\text{rad}}^{(6)}$ 1735.4057-1735.4057 67.5480-67.5480 1799.2044-1799.2044
$\Delta{E^{\prime}}_{\text{rel}}^{(6)}$ 129.500(20)-129.500(20) 5.485(21)-5.485(21) 134.689(20)-134.689(20)
$\Delta(E^{\prime}_{B}+E^{\prime}_{S})$ $153.998(26)$ $7.353(26)$ $160.864(26)$
$\Delta E_{\text{rad-rec}}^{(6)}$ $0.3105$ $0.0121$ $0.3219$
$\Delta{E^{\prime}}_{\text{rel-rec}}^{(6)}$ 0.674(11)-0.674(11) $9.303(11)$ 0.696(11)-0.696(11)
$\Delta E^{\prime}_{R}$ $0.5868(3)$ $0.0265(3)$ $0.6112(3)$
$\Delta E_{\text{total}}$ 1710.684(35)-1710.684(35) 56.339(35)-56.339(35) 1772.792(35)-1772.792(35)
[21] 1708.9(1)-1708.9(1)
The matrix elements for the operators that appear in the $m\alpha^{6}$-order correction $E^{(6m)}$, see Eq. (56), are calculated using the Hylleraas variational expansion of the wave function:
$\phi_{ijk}^{l_{1}l_{2}L}=r_{1}^{i}r_{2}^{j}r_{12}^{k}e^{-\alpha{r}_{1}-\beta{r}_{2}}Y^{l_{1}l_{2}}_{LM}(\hat{r}_{1},\hat{r}_{2})\,.$ (57)
In the above, $i,j,k$ are integers satisfying $i+j+k\leq\Omega$ with $\Omega$ being a parameter defining the size of the basis set, $\alpha$ and $\beta$ are nonlinear parameters, and $Y^{l_{1}l_{2}}_{LM}(\hat{r}_{1},\hat{r}_{2})$ is the vector coupled production of spherical harmonics. More details on the construction of the basis set for HD+ may be found in Refs. [8, 9, 33]. Most integrals can be calculated analytically using Perkins’ expansion for $r_{12}^{k}$ [34]. The procedure for handling singular integrals that appear in the evaluation can be found in Ref. [35, 36]. The divergent matrix elements $\langle V_{a}^{3}\rangle$, $\langle\bm{\varepsilon}_{a}^{2}\rangle$ and $\langle V_{a}\rho_{a}\rangle$ are calculated in the regularized form as defined in Eq. (70).
Table 1 shows the values of regularized divergent matrix elements that appear in the modified effective Hamiltonian, while Table 2 contains the numerical values for the finite matrix elements in Eq. (56), including the nonrelativistic energy $E_{0}$, the potential $U$, Eq. (28), the Breit Hamiltonians $H_{B}$ and $H_{R}$, Eqs. (6) and (7).
Table 3 presents the expectation values of the modified effective Hamiltonians that appear in the nonrecoil and recoil relativistic corrections ${E^{\prime}}_{\text{rel}}^{(6)}$ and ${E^{\prime}}_{\text{rel-rec}}^{(6)}$. Because of the regularization and transformation applied to the relativistic corrections, these modified effective Hamiltonians are derived by neglecting recoil corrections beyond the leading orders. Consequently, our theoretical framework exhibits a relative uncertainty of $\mathcal{O}((m_{e}/m_{p})^{2})$ for effective Hamiltonians $\delta H^{\prime}_{1}$, $\delta H^{\prime}_{2}$, $\delta H^{\prime}_{3}$, $\delta H^{\prime}_{4}$ and ${H^{\prime}}_{B}^{\text{1st}}$, even though our numerical precision is significantly higher.
[…]
The $m\alpha^{6}$ order corrections to transition frequencies in the HD+ ion are presented in Table 4. The uncertainty in the nonrecoil relativistic correction $\Delta{E^{\prime}}_{\text{rel}}^{(6)}$ and the recoil relativistic correction $\Delta{E^{\prime}}_{\text{rel-rec}}^{(6)}$ are estimated to be 20 Hz and 11 Hz respectively, primarily due to neglect of the higher-order recoil effects. The uncertainty due to the second-order contribution $\Delta(E^{\prime}_{B}+E_{S})$ is about 26 Hz, which is estimated by the last significant digit of the corresponding values of $E^{\prime}_{B}$ and $E_{S}$. The resulting $m\alpha^{6}$-order correction to the $(0,0)\!\to\!(0,1)$ transition is determined to be $1710.684(35)~\text{kHz}$, the uncertainty of which is three orders of magnitude smaller than the value reported in previous theoretical studies [21, 23]. As is seen, there is a discrepancy of approximately 1.8 kHz between the obtained value and the previous theory. We attribute this to the limitations of the adiabatic Born-Oppenheimer approximation employed in previous works.
In conclusion, we present numerical calculations for the first-order corrections at $m\alpha^{6}$- and $m\alpha^{6}(m/M)$-order in the three-body formalism for the hydrogen molecular ions. It relies heavily on the coordinate cutoff regularization. The regularization procedure is presented in detail, especially for the recoil corrections. Matrix elements appearing in the regularized effective Hamiltonians are calculated to high precision based on the Hylleraas basis set. The expectation values of effective Hamiltonians are presented with a relative uncertainty of $\mathcal{O}((m_{e}/m_{p})^{2})$. In combination with the numerical results for the second-order corrections [27], the total contribution of the $m\alpha^{6}$-order to the fundamental transitions in HD+ is obtained with precision of $35~\text{Hz}$, which is three orders of magnitude better than in the previous work. As stated in [21], the largest contribution to the uncertainty comes from the $m\alpha^{8}$ order correction, which is four times larger than the uncertainty in the $m\alpha^{6}$ order correction. Nevertheless, the results obtained in this work allow us to state that in theoretical studies of the spectra of vibrational transitions in molecular hydrogen ions, a relative accuracy of $10^{-12}$ or higher can be achieved.
Acknowledgements.
This research was supported by the National Natural Science Foundation of China (NSFC) under Grants Nos. 12393821. ZXZ acknowledges support from the Innovation Fund for Scientific and Technological Personnel of Hainna Province (Grant No. 225RC636). Chun Li was supported in part by the National Natural Science Foundation of China Grant (No. 11871271). T. Y. Shi was supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant Nos.XDB0920100 and XDB0920101), the National Natural Science Foundation of China (Grant No.12274423). Authors would like to thank helpful comments from Z.-C. Yan of UNB and J.-Ph. Karr of LKB.
Appendix A Coordinate cutoff regularization
In the coordinate cutoff regularization, the divergent integrals can be integrated by setting a lower limit for integration over the radial variable to some small parameter $r_{0}$. The finite distribution for $1/r^{3}$ integral was introduced by Araki [38] and Sucher [39] in connection with the $m\alpha^{5}$ order correction for the helium atom. The general analysis of the cutoff regularization for the $m\alpha^{6}$ order relativistic correcions has been given in Ref. [25], Section IV.A.
Here we present a brief summary of the main features of the cutoff regularization for bound states. In our case, all singularities are separated out in the form of singular integrals: $\langle V^{3}\rangle$, $\langle\bm{\varepsilon}^{2}\rangle$, $\langle V\imath\bm{\varepsilon}\mathbf{p}_{e}\rangle$ and $\langle V\rho\rangle$. The divergent part of those integrals may be expressed as follows:
[…]
where $\gamma_{E}$ is the Euler gamma constant and
[…]
In the above formulas, $V=z_{e}z_{N}/r$ is the Coulomb interaction between the electron and the nucleus, where $z_{e}$ and $z_{N}$ are the charges of the electron and the nucleus, respectively.
In the coordinate cutoff regularization, the finite integrals for these matrix elements are defined as follows:
$\begin{array}[]{@{}l}\displaystyle\langle V\rangle_{r}=(z_{e}z_{N})^{3}\lim_{r_{0}\rightarrow 0}\left\{\left\langle\frac{1}{r}\right\rangle_{r_{0}}+(\ln r_{0}+\gamma_{E})\langle 4\pi\delta(\bm{r})\rangle\right\},\\[8.53581pt] \displaystyle\langle\bm{\varepsilon}^{2}\rangle_{r}=(z_{e}z_{N})^{2}\lim_{r_{0}\rightarrow 0}\Bigg\{\left\langle\frac{1}{r}\right\rangle_{r_{0}}-\bigg[\frac{\langle 4\pi\delta(\bm{r})\rangle}{r_{0}}\\[8.53581pt] \displaystyle\hskip 99.58464pt+(\ln r_{0}+\gamma_{E})\langle 4\pi\delta^{\prime}(\bm{r})\rangle\bigg]\Bigg\},\\[8.53581pt] \displaystyle\langle V\rho\rangle_{r}=-(z_{e}z_{N})^{2}\langle 4\pi\delta^{\prime}(\mathbf{r})\rangle\,.\end{array}$ (70)
The numerical evaluation of the divergent integrals in Eq. (70) for the Hylleraas variational basis functions used in our work is discussed in Ref. [40].
Appendix B Divergent matrix elements
In order to express $H^{(6)}_{\rm rel}$ in Eq. (9) in terms of finite matrix elements, we transform the effective Hamiltonian and the second-order corrections into a set of regular and divergent (but redefined in the finite form) matrix elements, such as $\langle V_{a}\mathbf{p}_{e}^{2}V_{a}\rangle$, $\{\mathbf{p}_{e}^{4},V_{a}\}$, and $\langle\mathbf{p}_{e}^{2}V_{a}\mathbf{p}_{e}^{2}\rangle$, etc. Many of these matrix elements were already considered in Ref. [24]; here we write out exlicitly our previously obtained results and derive additional new expressions for matrix elements that appear in this work.
B.1 Matrix elements $\langle V_{a}^{2}\mathbf{p}_{e}^{2}\rangle$, $\langle V_{a}^{2}\mathbf{p}_{a}^{2}\rangle$, $\langle V_{a}\mathbf{p}_{e}^{2}V_{a}\rangle$, $\langle V_{a}\mathbf{p}_{a}^{2}V_{a}\rangle$ and $\langle{V}\mathbf{p}_{a}^{2}V\rangle$
In this section we will frequently use the following expression for $\mathbf{p}_{e}^{2}\psi_{0}$:
$\mathbf{p}_{e}^{2}\psi_{0}=\left[2m_{e}(E_{0}-V)-\sum_{a}\frac{m_{e}}{m_{a}}\mathbf{p}_{a}^{2}\right]\psi_{0}\,.$ (71)
which is equivalent to the Schrödinger equation (2).
The finite (regularized) distribution for the matrix element $\langle V_{a}^{2}\mathbf{p}_{e}^{2}\rangle$ can be determined as follows
[…]
$\langle V_{a}^{2}\mathbf{p}_{e}^{2}\rangle^{(m)}=-2\mu_{a}\langle V_{a}^{3}\rangle_{r}+\langle V_{a}^{2}\mathbf{p}_{e}^{2}\rangle^{\prime}\,$, (75)
where the first term on the right-hand side is regularized divergent, and the second term is finite
$\langle V_{a}^{2}\mathbf{p}_{e}^{2}\rangle^{\prime}$ $=\mu_{a}\bigg\{2\left[E_{0}\langle V_{a}^{2}\rangle-\langle V_{a}^{2}(V_{b}+V_{12})\rangle\right]$
$\qquad-\frac{1}{m_{a}}\langle V_{a}^{2}(\mathbf{p}_{a}^{2}-\mathbf{p}_{e}^{2})\rangle-\frac{1}{m_{b}}\langle V_{a}^{2}\mathbf{p}_{b}^{2}\rangle\bigg\}\,.$ (76)
In $\langle V_{a}^{2}\mathbf{p}_{e}^{2}\rangle^{\prime}$, indexes $a$ and $b$ stand for the two nuclei with $b\neq{a}$, and $\mu_{a}=m_{a}m_{e}/(m_{a}+m_{e})$ are the reduced masses.
The finite matrix element for $\langle V_{a}^{2}\mathbf{p}_{a}^{2}\rangle$ then can be determined
$\langle V_{a}^{2}\mathbf{p}_{a}^{2}\rangle^{(m)}=-2\mu_{a}\langle V_{a}^{3}\rangle_{r}+\langle V_{a}^{2}\mathbf{p}_{a}^{2}\rangle^{\prime}\,$, (77)
where
$\langle V_{a}^{2}\mathbf{p}_{a}^{2}\rangle^{\prime}=\langle V_{a}^{2}\mathbf{p}_{e}^{2}\rangle^{\prime}+\langle V_{a}^{2}(\mathbf{p}_{a}^{2}-\mathbf{p}_{e}^{2})\rangle\,.$ (78)
The matrix element $\langle V_{a}\mathbf{p}_{e}^{2}V_{a}\rangle$ may be expressed as follows (see Eq. (A6) in Ref. [24])
$\langle V_{a}\mathbf{p}_{e}^{2}V_{a}\rangle^{(m)}=\langle\bm{\varepsilon}_{a}^{2}\rangle_{r}-2\mu_{a}\langle V_{a}^{3}\rangle_{r}+\langle V_{a}^{2}\mathbf{p}_{e}^{2}\rangle^{\prime}\,.$ (79)
The matrix element $\langle V_{a}\mathbf{p}_{a}^{2}V_{a}\rangle$ is expressed in the similar form
$\langle V_{a}\mathbf{p}_{a}^{2}V_{a}\rangle^{(m)}=\langle\bm{\varepsilon}_{a}^{2}\rangle_{r}-2\mu_{a}\langle V_{a}^{3}\rangle_{r}+\langle V_{a}^{2}\mathbf{p}_{a}^{2}\rangle^{\prime}\,.$ (80)
And the cross-term
$\langle V_{a}\mathbf{p}_{e}^{2}V_{b}\rangle=\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{b}\rangle+\langle V_{a}V_{b}\mathbf{p}_{e}^{2}\rangle\,.$ (81)
In the latter case all the operators are finite.
Now we may get the matrix element $\langle V\mathbf{p}_{a}^{2}V\rangle$:
$\langle V\mathbf{p}_{a}^{2}V\rangle^{(m)}$ $=\langle V_{a}\mathbf{p}_{a}^{2}V_{a}\rangle^{(m)}+\langle\bm{\varepsilon}_{12}^{2}\rangle_{r}\pm 2\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{12}\rangle+\langle(V^{2}-V_{a}^{2})\mathbf{p}_{a}^{2}\rangle\,.$ (82)
$\langle V\mathbf{p}_{a}^{2}V\rangle^{(m)}=\langle\bm{\varepsilon}_{12}^{2}\rangle_{r}+\langle\bm{\varepsilon}_{a}^{2}\rangle_{r}-2\mu_{a}\langle V_{a}^{3}\rangle_{r}+\langle V\mathbf{p}_{a}^{2}V\rangle^{\prime}\,$, (83)
where
$\langle V\mathbf{p}_{a}^{2}V\rangle^{\prime}=\langle{V}_{a}^{2}\mathbf{p}_{a}^{2}\rangle^{\prime}\pm 2\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{12}\rangle+\langle(V^{2}-V_{a}^{2})\mathbf{p}_{a}^{2}\rangle\,.$ (84)
In the above derivation, singular integrals involving $r_{12}$ appear implicitly in the form $\langle\bm{\varepsilon}_{12}^{2}\rangle+\langle V_{12}^{2}\mathbf{p}_{a}^{2}\rangle$, which can be transferred as (see Eq. (21) of Ref. [25])
$\langle\bm{\varepsilon}_{12}^{2}\rangle+\langle V_{12}^{2}\mathbf{p}_{a}^{2}\rangle=\langle V_{12}\rho_{12}\rangle+\langle\mathbf{p}_{a}V_{12}^{2}\mathbf{p}_{a}\rangle\,.$
The first term of the right-hand side is sigular, while the second term is regular. In case of the molecular ions, the distribution $\rho_{12}$ is vanishingly small. As a consequence, the nucleus-nucleus divergence in $\langle V\mathbf{p}_{a}^{2}V\rangle$ cancels out algebraically. Therefore, the matrix element $\langle V\mathbf{p}_{a}^{2}V\rangle$ can be evaluated using the formula above within the cut-off regularization scheme.
All the modified matrix elements written on the right-hand side of equations in this section are finite and may be used in numerical calculations.
Yet another relation which connects the three basic singular distributions in the three-body case should be written
[…]
$\langle V_{a}\rho_{a}\rangle-\langle\bm{\varepsilon}_{a}^{2}\rangle+2\mu_{a}\langle V_{a}^{3}\rangle=\langle V_{a}\rho_{a}\rangle^{\prime}\,$, (86)
where $\langle V_{a}\rho_{a}\rangle^{\prime}$ is defined as follows
$\langle V_{a}\rho_{a}\rangle^{\prime}=\langle V_{a}^{2}\mathbf{p}_{e}^{2}\rangle^{\prime}-\langle\mathbf{p}_{e}V_{a}^{2}\mathbf{p}_{e}\rangle\,.$ (87)
B.2 Matrix elements $\langle\{\mathbf{p}_{e}^{2},\rho_{a}\}\rangle$, $\langle\{\mathbf{p}_{e}^{4},V_{a}\}\rangle$, $\langle\{\mathbf{p}_{e}^{4},V_{12}\}\rangle$ and $\langle\mathbf{p}_{e}^{2}V_{a}\mathbf{p}_{e}^{2}\rangle$
The finite part of the matrix element $\langle\{\mathbf{p}_{e}^{2},\rho_{a}\}\rangle$ can be expressed using Eqs. (71) and (86):
$\begin{array}[]{@{}l}\displaystyle\langle\{\mathbf{p}_{e}^{2},\rho_{a}\}\rangle^{(m)}=-4\mu_{a}\langle V_{a}\rho_{a}\rangle_{r}+\mu_{a}\bigg\{4\left[E_{0}\langle\rho_{a}\rangle-\langle(V_{b}\!+\!V_{12})\rho_{a}\rangle\right]\\[8.53581pt] \displaystyle\hskip 150.79959pt-\frac{2}{m_{a}}\langle(\mathbf{p}_{a}^{2}\!-\!\mathbf{p}_{e}^{2})\rho_{a}\rangle-\frac{2}{m_{b}}\langle\mathbf{p}_{b}^{2}\rho_{a}\rangle\bigg\}\,.\end{array}$ (90)
$\langle\{\mathbf{p}_{e}^{2},\rho_{a}\}\rangle^{(m)}=-4\mu_{a}\left(\langle\bm{\varepsilon}_{a}^{2}\rangle_{r}-2\mu_{a}\langle{V}_{a}^{3}\rangle_{r}\right)+\langle\{\mathbf{p}_{e}^{2},\rho_{a}\}\rangle^{\prime}\,$, (91)
where
$\langle\{\mathbf{p}_{e}^{2},\rho_{a}\}\rangle^{\prime}$ $=-4\mu_{a}\langle V_{a}\rho_{a}\rangle^{\prime}+\mu_{a}\bigg\{4\left[E_{0}\langle\rho_{a}\rangle-\langle(V_{b}\!+\!V_{12})\rho_{a}\rangle\right]$
$-\frac{2}{m_{a}}\langle(\mathbf{p}_{a}^{2}\!-\!\mathbf{p}_{e}^{2})\rho_{a}\rangle-\frac{2}{m_{b}}\langle\mathbf{p}_{b}^{2}\rho_{a}\rangle\bigg\}\,.$ (92)
Next operator:
$\langle\{\mathbf{p}_{e}^{4},V_{a}\}\rangle^{(m)}$ $=-4m_{e}\left(1\!-\!\frac{m_{e}}{m_{a}}\right)\left(\langle\bm{\varepsilon}_{a}^{2}\rangle_{r}-2\mu_{a}\langle{V}_{a}^{3}\rangle_{r}\right)$
$\qquad+\langle\{\mathbf{p}_{e}^{4},V_{a}\}\rangle^{\prime}\,$, (93)
where
$\begin{array}[]{@{}l}\displaystyle\langle\{\mathbf{p}_{e}^{4},V_{a}\}\rangle^{\prime}=-4m_{e}\left(1\!-\!\frac{m_{e}}{m_{a}}\right)\langle V_{a}\mathbf{p}_{e}^{2}V_{a}\rangle^{\prime}\\[8.53581pt] \displaystyle\hskip 56.9055pt+4m_{e}E_{0}\left\langle V_{a}\left(\mathbf{p}_{e}^{2}-\frac{m_{e}}{m_{a}}\mathbf{p}_{a}^{2}-\frac{m_{e}}{m_{b}}\mathbf{p}_{b}^{2}\right)\right\rangle\\[8.53581pt] \displaystyle\hskip 56.9055pt-4m_{e}\left[\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{b}\rangle+\langle V_{a}(V_{b}+V_{12})\mathbf{p}_{e}^{2}\rangle\right]\\[8.53581pt] \displaystyle\hskip 56.9055pt+\frac{4m_{e}^{2}}{m_{a}}\Bigl[\pm\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{12}\rangle+\langle V_{a}^{2}(\mathbf{p}_{a}^{2}-\mathbf{p}_{e}^{2})\rangle\\[8.53581pt] \displaystyle\hskip 56.9055pt+\langle V_{a}(V_{b}+V_{12})\mathbf{p}_{a}^{2}\rangle\Bigr]+\frac{4m_{e}^{2}}{m_{b}}\langle V_{a}V\mathbf{p}_{b}^{2}\rangle\,.\end{array}$ (99)
Here higher-order terms in the expansion in $(m_{e}/m_{a})^{n}$ are ignored. All matrix elements are regular.
The matrix element $\langle\mathbf{p}_{e}^{2}V_{a}\mathbf{p}_{e}^{2}\rangle$ can be expanded similarly:
$\langle\mathbf{p}_{e}^{2}V_{a}\mathbf{p}_{e}^{2}\rangle^{(m)}=4\mu_{a}\left(1\!-\!\frac{m_{e}}{m_{a}}\right)\langle V_{a}^{3}\rangle_{r}+\langle\mathbf{p}_{e}^{2}V_{a}\mathbf{p}_{e}^{2}\rangle^{\prime}\,$, (100)
where
$\begin{array}[]{@{}l}\displaystyle\langle\mathbf{p}_{e}^{2}V_{a}\mathbf{p}_{e}^{2}\rangle^{\prime}=-2m_{e}\left(1\!-\!\frac{1}{m_{a}}\right)\langle V_{a}^{2}\mathbf{p}_{e}^{2}\rangle^{\prime}\\[8.53581pt] \displaystyle\hskip 28.45274pt\!+\!2m_{e}E_{0}\left\langle V_{a}\left(\mathbf{p}_{e}^{2}\!-\!\frac{m_{e}}{m_{a}}\mathbf{p}_{a}^{2}\!-\!\frac{m_{e}}{m_{b}}\mathbf{p}_{b}^{2}\right)\right\rangle\!+\!\frac{2m_{e}^{2}}{m_{b}}\langle V_{a}V\mathbf{p}_{b}^{2}\rangle\hskip-28.45274pt\\[8.53581pt] \displaystyle\hskip 28.45274pt-2m_{e}\left[\langle V_{a}(V_{b}\!+\!V_{12})\mathbf{p}_{e}^{2}\rangle-\frac{m_{e}}{m_{a}}\langle V_{a}(V_{b}\!+\!V_{12})\mathbf{p}_{a}^{2}\rangle\right]\\[8.53581pt] \displaystyle\hskip 28.45274pt+\frac{2m_{e}^{2}}{m_{a}}\langle V_{a}^{2}(\mathbf{p}_{a}^{2}-\mathbf{p}_{e}^{2})\rangle.\end{array}$ (105)
It can be seen that the matrix elements $\langle\{\mathbf{p}_{e}^{4},V_{a}\}\rangle^{(m)}$ and $\langle\mathbf{p}_{e}^{2}V_{a}\mathbf{p}_{e}^{2}\rangle^{(m)}$ are related by the following identity
$\langle\{\mathbf{p}_{e}^{4},V_{a}\}\rangle^{(m)}-2\langle\mathbf{p}_{e}^{2}V_{a}\mathbf{p}_{e}^{2}\rangle^{(m)}=4m_{e}\left(1\!-\!\frac{m_{e}}{m_{a}}\right)\langle\bm{\varepsilon}_{a}^{2}\rangle_{r}$
$\qquad-4m_{e}\bigg[\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{b}\rangle\mp\frac{m_{e}}{m_{a}}\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{12}\rangle\bigg]$ (106)
The regular matrix element $\langle\{\mathbf{p}_{e}^{4},V_{12}\}\rangle$ can be calculated as follows:
$\langle\{\mathbf{p}_{e}^{4},V_{12}\}\rangle=2\langle\mathbf{p}_{e}^{2}V_{12}\mathbf{p}_{e}^{2}\rangle\,.$ (107)
B.3 Matrix element $\langle U(E_{0}-H_{0})U\rangle$
The matrix element $\langle U(E_{0}-H_{0})U\rangle$ is expanded by inserting Eq. (28):
$\langle{U}(E_{0}\!-\!H_{0})U\rangle=\lambda_{1}^{2}\langle{V}_{1}(E_{0}\!-\!H_{0})V_{1}\rangle+\lambda_{2}^{2}\langle{V}_{2}(E_{0}\!-\!H_{0})V_{2}\rangle$
$\qquad+2\lambda_{1}\lambda_{2}\langle{V}_{1}(E_{0}\!-\!H_{0})V_{2}\rangle\,$, (108)
where singular terms of the type $\langle V_{a}(E_{0}-H_{0})V_{a}\rangle$ can be simplified:
$\langle{V}_{a}(E_{0}-H_{0})V_{a}\rangle$ $=\langle{V}_{a}^{2}E_{0}\rangle-\langle{V}_{a}^{2}V\rangle-\frac{1}{2m_{a}}\langle{V}_{a}\mathbf{p}_{a}^{2}V_{a}\rangle$
$\quad-\frac{1}{2m_{b}}\langle{V}_{a}\mathbf{p}_{b}^{2}V_{a}\rangle-\frac{1}{2m_{e}}\langle{V}_{a}\mathbf{p}_{e}^{2}V_{a}\rangle$
$=-\frac{1}{2m_{e}}\left(1\!+\!\frac{m_{e}}{m_{a}}\right)\langle\bm{\varepsilon}_{a}^{2}\rangle_{r}\,.$ (109)
As was shown, the singular terms $\langle\bm{\varepsilon}_{a}^{2}\rangle$ after summation within the nonrecoil contribution are cancelled out.
The matrix element $\langle V_{1}(E_{0}\!-\!H_{0})V_{2}\rangle$ may be simplified by substituting the Hamiltonian (2):
$\langle V_{1}(E_{0}-H_{0})V_{2}\rangle$ $=\left\langle V_{1}\left(E_{0}-\frac{\mathbf{p}_{e}^{2}}{2m_{e}}-\frac{\mathbf{p}_{1}^{2}}{2m_{1}}-\frac{\mathbf{p}_{2}^{2}}{2m_{2}}-V\right)V_{2}\right\rangle$
$=E_{0}\langle V_{1}V_{2}\rangle-\frac{1}{2m_{e}}\langle V_{1}\mathbf{p}_{e}^{2}V_{2}\rangle-\frac{1}{2m_{1}}\langle V_{1}V_{2}\mathbf{p}_{1}^{2}\rangle$
$\quad-\frac{1}{2m_{2}}\langle V_{1}V_{2}\mathbf{p}_{2}^{2}\rangle-\langle V_{1}V_{2}V\rangle$
$=E_{0}\langle V_{1}V_{2}\rangle-\frac{1}{2m_{e}}\langle\bm{\varepsilon}_{1}\bm{\varepsilon}_{2}\rangle-\left\langle V_{1}V_{2}H_{0}\right\rangle$
$=-\frac{1}{2m_{e}}\langle\bm{\varepsilon}_{1}\bm{\varepsilon}_{2}\rangle\,.$ (110)
The above derivation uses the identity (81) and the fact that the wave function is a solution of the Schrödinger equation $(E_{0}\!-\!H_{0})\psi_{0}=0$.
Appendix C Derivation of modified effective Hamiltonian
In this appendix, the singularity separation for the effective Hamiltonian $\delta H_{1}$-$\delta H_{6}$ is performed in detail, along with the second-order contribution $E_{B}$ and $E_{R}$.
It has been proved in many articles that the singularities of the $m\alpha^{6}$-order nonrecoil contributions are algebraically cancelled. Therefore, the regular parts of those divergent matrix elements are removed as well. On the other hand, the singularities of the $m\alpha^{6}(m/M)$-order recoil contributions are removed while its regular part should be kept in the final expression.
C.1 expansion of $\delta H_{1}$
Let us consider the effective Hamiltonian $\delta H_{1}$,
$\delta H_{1}=\frac{\mathbf{p}_{e}^{6}}{16m_{e}^{5}}\,.$ (111)
Using Eq. (71), the matrix element $\langle\mathbf{p}_{e}^{6}\rangle$ can be expanded as follows:
$\langle\mathbf{p}_{e}^{6}\rangle^{(m)}$ $=2m_{e}E_{0}\langle\mathbf{p}_{e}^{4}\rangle-m_{e}\langle\{\mathbf{p}_{e}^{4},V_{12}\}\rangle$
$\quad-\sum_{a}m_{e}\left(\langle\{\mathbf{p}_{e}^{4},V_{a}\}\rangle^{(m)}+\frac{1}{2m_{a}}\langle\{\mathbf{p}_{e}^{4},\mathbf{p}_{a}^{2}\}\rangle^{(m)}\right)\,$, (112)
where the last term has the expansion as follows (ignoring higher-order recoil corrections)
$\frac{1}{m_{a}}\langle\{\mathbf{p}_{e}^{4},\mathbf{p}_{a}^{2}\}\rangle^{(m)}$ $=\frac{8m_{e}^{2}}{m_{a}}\bigg(E_{0}^{2}\langle\mathbf{p}_{a}^{2}\rangle-2E_{0}\langle V\mathbf{p}_{a}^{2}\rangle$
$\quad+\langle V\mathbf{p}_{a}^{2}V\rangle^{(m)}\bigg)+O(1/m_{a,b}^{2})\,.$ (113)
The terms $\langle\{\mathbf{p}_{e}^{4},V_{a}\}\rangle^{(m)}$ and $\langle V\mathbf{p}_{a}^{2}V\rangle^{(m)}$ can be taken from Eqs. (B.2) and (83) respectively. The regular matrix elements $\langle\{\mathbf{p}_{e}^{4},V_{12}\}\rangle$ can be calculated directly, see Eq. (107).
Inserting expressions of all matrix elements into $\langle\mathbf{p}^{6}_{e}\rangle$, we can obtain the explicit expansion for the modified effective Hamiltonian $\langle\delta H_{1}^{(m)}\rangle$:
$\begin{array}[]{@{}l}\displaystyle\langle\delta H^{\prime}_{1}\rangle=\frac{1}{16m_{e}^{4}}\Bigg\{2E_{0}\langle\mathbf{p}_{e}^{4}\rangle-\langle\{\mathbf{p}_{e}^{4},V_{12}\}\rangle\\[8.53581pt] \displaystyle\hskip 56.9055pt-\sum_{a}\bigg[\langle\{\mathbf{p}_{e}^{4},V_{a}\}\rangle^{\prime}\\[8.53581pt] \displaystyle\hskip 56.9055pt+\frac{4m_{e}^{2}}{m_{a}}\left(E_{0}^{2}\langle\mathbf{p}_{a}^{2}\rangle-2E_{0}\langle V\mathbf{p}_{a}^{2}\rangle+\langle V\mathbf{p}_{a}^{2}V\rangle^{\prime}\right)\bigg]\Bigg\}\,.\end{array}$ (117)
The singular part corresponding to $\langle\delta H_{1}\rangle$ takes the form
$S_{1}=\sum_{a}\frac{1}{4m_{e}^{3}}\bigg[\left(1-\frac{2m_{e}}{m_{a}}\right)\langle\bm{\varepsilon}_{a}^{2}\rangle-2m_{e}\left(1-\frac{3m_{e}}{m_{a}}\right)\langle V_{a}^{3}\rangle\bigg]\,.$ (118)
C.2 expansion of $\delta H_{2}$
The next term of the effective Hamiltonian is $\delta H_{2}$,
$\delta H_{2}=\frac{1}{128m_{e}^{4}}[\mathbf{p}_{e}^{2},[\mathbf{p}_{e}^{2},V]]+\sum_{a}\frac{3}{64m_{e}^{4}}\{\mathbf{p}_{e}^{2},\rho_{a}\}\,.$ (119)
The expectation value for the first operator can be simplified as follows
$\langle[\mathbf{p}_{e}^{2},[\mathbf{p}_{e}^{2},V]]\rangle^{(m)}$ $=\langle\{\mathbf{p}_{e}^{4},V\}\rangle^{(m)}-2\langle\mathbf{p}_{e}^{2}V\mathbf{p}_{e}^{2}\rangle^{(m)}$
$=\sum_{a}\bigg(\langle\{\mathbf{p}_{e}^{4},V_{a}\}\rangle^{(m)}-2\langle\mathbf{p}_{e}^{2}V_{a}\mathbf{p}_{e}^{2}\rangle^{(m)}\bigg)$
$=-4m_{e}\sum_{a}\bigg[\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{b}\rangle\mp\frac{m_{e}}{m_{a}}\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{12}\rangle$
$\quad+\left(1-\frac{m_{e}}{m_{a}}\right)\langle\bm{\varepsilon}_{a}^{2}\rangle_{r}\bigg]\,.$ (120)
Here the identity of Eq. (B.2) is adopted to get the final result.
The second term of $\delta H_{2}$ is proportional to $\{\mathbf{p}_{e}^{2},\rho_{a}\}$ and can be treated as in Eq. (91).
Combining the above matrix elements, one can obtained the modified $\delta H_{2}$ as
$\langle\delta H^{\prime}_{2}\rangle=\sum_{a}\bigg[-\frac{1}{32m_{e}^{3}}\bigg(\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{b}\rangle\mp\frac{m_{e}}{m_{a}}\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{12}\rangle\bigg)+\frac{3}{64m_{e}^{4}}\langle\{p_{e}^{2},\rho_{a}\}\rangle^{\prime}\bigg]\,$, (121)
And the singular part is
$S_{2}=\sum_{a}\bigg[-\frac{7}{32m_{e}^{3}}\left(1-\frac{m_{e}}{m_{a}}\right)\langle\bm{\varepsilon}_{a}^{2}\rangle+\frac{3}{8m_{e}^{2}}\left(1-\frac{2m_{e}}{m_{a}}\right)\langle V_{a}^{3}\rangle\bigg]\,.$ (122)
C.3 expansion of $\delta H_{3}$
The effective Hamiltonian $\delta H_{3}$ has the form of recoil contribution, but is identified as part the nonrecoil contribution,
$\delta H_{3}=\sum_{a}-\frac{1}{32m_{e}^{3}m_{a}}[\mathbf{p}_{e}^{2},[\mathbf{p}_{a}^{2},V_{a}]]\,.$ (123)
The matrix element $\langle[\mathbf{p}_{e}^{2},[\mathbf{p}_{a}^{2},V_{a}]]\rangle/m_{a}$ is transformed in a similar way as in Eq. (C.2):
$\frac{1}{m_{a}}\langle[\mathbf{p}_{e}^{2},[\mathbf{p}_{a}^{2},V_{a}]]\rangle^{(m)}$ $=\frac{2m_{e}}{m_{a}}\langle[[\mathbf{p}_{a}^{2},V_{a}],V]\rangle^{(m)}$
$=\frac{4m_{e}}{m_{a}}\left(\langle VV_{a}\mathbf{p}_{a}^{2}\rangle^{(m)}-\langle V_{a}\mathbf{p}_{a}^{2}V\rangle^{(m)}\right)$
$=\frac{4m_{e}}{m_{a}}\bigg(\langle V_{a}^{2}\mathbf{p}_{a}^{2}\rangle^{(m)}-\langle V_{a}\mathbf{p}_{a}^{2}V_{a}\rangle^{(m)}$
$\quad+\langle V_{a}V_{12}\mathbf{p}_{a}^{2}\rangle-\langle V_{a}\mathbf{p}_{a}^{2}V_{12}\rangle\bigg)$
$=-\frac{4m_{e}}{m_{a}}\left(\langle\bm{\varepsilon}_{a}^{2}\rangle_{r}\pm\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{12}\rangle\right)\,.$ (124)
Therefore, one gets the modified $\delta H_{3}$ in the form
$\langle\delta H^{\prime}_{3}\rangle=\sum_{a}\mp\frac{1}{8m_{e}^{2}m_{a}}\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{12}\rangle\,$, (125)
while the singular part is separated as
$S_{3}=\sum_{a}\frac{1}{8m_{e}^{2}m_{a}}\langle\bm{\varepsilon}_{a}^{2}\rangle\,.$ (126)
C.4 expansion of $\delta H_{4}$
Now we turn to the effective Hamiltonian $\delta H_{4}$, which has been transferred as Eq. (III.2):
$\langle\delta H_{4}\rangle=-\frac{E_{0}}{m_{e}}\langle H_{R}\rangle-\sum_{a}\frac{z_{e}z_{a}}{2m_{e}^{2}m_{a}}\langle\{V,~p_{e}^{i}\mathcal{W}^{ij}_{a}p_{a}^{j}\}\rangle+O(1/m_{a,b}^{2})\,.$ (127)
the divergent matrix element $\langle\{V,~p_{e}^{i}\mathcal{W}^{ij}_{a}p_{a}^{j}\}\rangle$ is expanded as Eq. (III.2) by acting momentum operators on the Coulomb potential,
$\langle\{V,p_{e}^{i}\mathcal{W}^{ij}_{a}p_{a}^{j}\}\rangle$ $=2\langle p_{e}^{i}\mathcal{W}^{ij}_{a}Vp_{a}^{j}\rangle-\langle\imath\varepsilon_{a}^{i}\mathcal{W}_{a}^{ij}p_{a}^{j}\rangle-\langle p_{e}^{i}\mathcal{W}_{a}^{ij}\imath\varepsilon_{a}^{j}\rangle\,.$ (128)
Let us consider the second term in Eq. (128). Adopting the identity $\mathbf{p}_{e}+\mathbf{p}_{1}+\mathbf{p}_{2}=0$, it can be split into two terms:
$z_{e}z_{a}\langle\imath\varepsilon_{a}^{i}\mathcal{W}_{a}^{ij}p_{a}^{j}\rangle=\langle V_{a}\imath\bm{\varepsilon}_{a}\mathbf{p}_{a}\rangle=-\langle V_{a}\imath\bm{\varepsilon}_{a}\mathbf{p}_{e}\rangle-\langle V_{a}\imath\bm{\varepsilon}_{a}\mathbf{p}_{b}\rangle\,.$ (129)
One can see that the first term is singular and the second term is regular. Meanwhile, it has been numerical verfied that matrix elements $\langle V_{a}\imath\bm{\varepsilon}_{a}\mathbf{p}_{b}\rangle$ are small enough to be ignored. We find that the last matrix element in Eq. (128) is singular,
$z_{e}z_{a}\langle p_{e}^{i}\mathcal{W}^{ij}_{a}\imath\varepsilon_{a}^{j}\rangle=\langle\mathbf{p}_{e}\imath\bm{\varepsilon}_{a}V_{a}\rangle\,.$ (130)
Inserting all matrix elements into Eq. (III.2) yields the modified $\langle\delta H_{4}^{(m)}\rangle$ while divergent matrix elements are removed
$\langle\delta H^{\prime}_{4}\rangle=-\frac{E_{0}}{m_{e}}\langle H_{R}\rangle-\sum_{a}\frac{1}{2m_{e}^{2}m_{a}}z_{e}z_{a}\langle 2p_{e}^{i}\mathcal{W}^{ij}_{a}Vp_{a}^{j}\rangle\,$, (131)
and the singularity of $\delta H_{4}$ is then has a form:
$S_{4}=-\sum_{a}\frac{1}{2m_{e}^{2}m_{a}}\left(\langle V_{a}\rho_{a}\rangle+\langle\bm{\varepsilon}_{a}^{2}\rangle\right)\,.$ (132)
C.5 expansion of $\delta H_{5}$
The modified effective Hamiltonian $\langle\delta H_{5}^{(m)}\rangle$ can be obtained easily,
$\langle\delta H^{\prime}_{5}\rangle=\sum_{a}\frac{1}{8m_{e}^{2}m_{a}}\bigg[\langle\bm{p}_{e}V_{a}^{2}\bm{p}_{e}\rangle+3\langle(\bm{p}_{e}\cdot\hat{r}_{a})V_{a}^{2}(\hat{r}_{a}\cdot\bm{p}_{e})\rangle\bigg]\,$, (133)
while the singularity of this term is
$S_{5}=\sum_{a}\frac{1}{4m_{e}^{2}m_{a}}\langle\bm{\varepsilon}_{a}^{2}\rangle\,.$ (134)
C.6 expansion of $\delta H_{6}$
The effective Hamiltonian $\delta H_{6}$ has a formula of:
$\delta H_{6}=\sum_{a}-\frac{z_{e}z_{a}}{m_{e}m_{a}}\bigg\{[p_{e}^{i},V]\mathcal{X}^{ij}_{a}[V,p_{a}^{j}]+p_{e}^{i}\left[\mathcal{X}^{ij}_{a},\frac{\mathbf{p}_{e}^{2}}{2m_{e}}\right][V,p_{a}^{j}]\bigg\}\,.$ (135)
The first matrix element is singular and can be expanded by performing $p_{e}^{i},V]$ and $V,p_{a}^{j}]$:
$z_{e}z_{a}\langle[p_{e}^{i},V]\mathcal{X}^{ij}_{a}[V,p_{a}^{j}]\rangle\!=\!z_{e}z_{a}\bigg\langle\imath(\varepsilon_{a}\!+\!\varepsilon_{b})^{i}\mathcal{X}^{ij}_{a}\imath\varepsilon_{a}^{j}\bigg\rangle$
$\qquad\!\mp\!z_{e}z_{a}\bigg\langle\imath(\varepsilon_{a}\!+\!\varepsilon_{b})^{i}\mathcal{X}^{ij}_{a}\imath\bm{\varepsilon}_{12}^{j}\bigg\rangle$
$\quad=\frac{1}{4}\langle V_{a}^{3}\rangle+\frac{z_{e}z_{a}}{4}\langle r_{a}\bm{\varepsilon}_{a}\bm{\varepsilon}_{b}\rangle\mp z_{e}z_{a}\bigg\langle\imath(\varepsilon_{a}+\varepsilon_{b})^{i}\mathcal{X}^{ij}_{a}\imath\bm{\varepsilon}_{12}^{j}\bigg\rangle\,.$ (136)
Let us evaluate the second term of $\delta H_{6}$. Firstly, we act by the operator $\mathbf{p}_{e}^{2}$ on $\mathcal{X}^{ij}_{a}$ to expand the commutator
$\left[\mathcal{X}^{ij}_{a},\frac{\mathbf{p}_{e}^{2}}{2m_{e}}\right]$ $=\frac{1}{2m_{e}}\left[\nabla^{2},\mathcal{X}^{ij}_{a}\right]=\frac{1}{m_{e}}\bigg[\left(\partial^{k}\mathcal{X}^{ij}_{a}\right)\partial^{k}+\frac{1}{2}\left(\partial^{k}\partial^{k}\mathcal{X}^{ij}_{a}\right)\bigg]\,$, (137)
where
$\partial^{k}\mathcal{X}^{ij}=\frac{1}{8}\left(\delta^{ki}\frac{r^{j}}{r}+\delta^{kj}\frac{r^{i}}{r}-\frac{r^{i}r^{j}r^{k}}{r^{3}}-3\delta^{ij}\frac{r^{k}}{r}\right)\,$, (138)
and
$\partial^{k}\partial^{k}\mathcal{X}^{ij}=-\frac{1}{2r}\left(\delta^{ij}+\frac{r^{i}r^{j}}{r^{2}}\right)\,.$ (139)
After that, the second term of $\delta H_{6}$ can be split into two terms. The first term is
$z_{e}z_{a}\bigg\langle p_{e}^{i}\left(\partial_{e}^{k}\mathcal{X}^{ij}_{a}\right)\partial_{e}^{k}\imath\varepsilon_{a}^{j}\bigg\rangle=\imath(z_{e}z_{a})^{2}\bigg\langle p_{e}^{i}\left(\partial_{e}^{k}\mathcal{X}^{ij}_{a}\right)\bigg[\frac{\delta^{kj}}{3}4\pi\delta(\mathbf{r}_{a})$
$\quad+\frac{1}{r_{a}^{3}}\left(\delta^{kj}-3\frac{r_{a}^{j}r_{a}^{k}}{r_{a}^{2}}\right)\bigg]\bigg\rangle+z_{e}z_{a}\bigg\langle p_{e}\left(\partial_{e}^{k}\mathcal{X}^{ij}_{a}\right)\imath\varepsilon_{a}^{j}\partial_{e}^{k}\bigg\rangle$
$=\frac{3}{4}\langle\mathbf{p}_{e}\imath\bm{\varepsilon}_{a}V_{a}\rangle-\frac{1}{8}\bigg[\langle\mathbf{p}_{e}V_{a}^{2}\mathbf{p}_{e}\rangle-3\left\langle(\mathbf{p}_{e}\cdot\mathbf{r}_{a})\frac{V_{a}^{2}}{r_{a}^{2}}(\mathbf{r}_{a}\cdot\mathbf{p}_{e})\right\rangle\bigg]\,.$ (140)
The second one is
$\frac{z_{e}z_{a}}{2}\bigg\langle p_{e}^{i}(\partial_{e}^{k}\partial_{e}^{k}\mathcal{X}^{ij}_{a})\imath\varepsilon_{a}^{j}\bigg\rangle$ $=\frac{\imath(z_{e}z_{a})^{2}}{2}\bigg\langle p_{e}^{i}\left(-\frac{1}{2}\right)\frac{1}{r_{a}}\left(\delta^{ij}+\frac{r_{a}^{i}r_{a}^{j}}{r_{a}^{2}}\right)\frac{r_{a}^{j}}{r_{a}^{3}}\bigg\rangle$
$=-\frac{1}{2}\langle\mathbf{p}_{e}\imath\bm{\varepsilon}_{a}V_{a}\rangle\,.$ (141)
Thus, the second term of $\delta H_{6}$ can be evaluated as
$z_{e}z_{a}\bigg\langle p_{e}^{i}\left[\mathcal{X}^{ij}_{a},\frac{\mathbf{p}_{e}^{2}}{2m_{e}}\right][V,p_{a}^{j}]\bigg\rangle=-\frac{1}{8m_{e}}\bigg[\left\langle\mathbf{p}_{e}V_{a}^{2}\mathbf{p}_{e}\right\rangle$
$\qquad-3\left\langle(\mathbf{p}_{e}\cdot\mathbf{r}_{a})\!\frac{V_{a}^{2}}{r_{a}^{2}}\!(\mathbf{r}_{a}\cdot\mathbf{p}_{e})\right\rangle\bigg]-\frac{1}{8m_{e}}\left[\langle V_{a}\rho_{a}\rangle+\langle\bm{\varepsilon}_{a}^{2}\rangle\right]\,.$ (142)
Now we can insert Eqs. (C.6) and (C.6) into $\langle\delta H_{6}\rangle$ to obtained the expectation value of the modified effective Hamiltonian
$\langle\delta H^{\prime}_{6}\rangle$ $=\sum_{a}\frac{1}{m_{e}m_{a}}\bigg\{\frac{1}{8m_{e}}\bigg[\left\langle\mathbf{p}_{e}V_{a}^{2}\mathbf{p}_{e}\right\rangle$
$\quad-3\left\langle(\mathbf{p}_{e}\cdot\mathbf{r}_{a})\!\frac{V_{a}^{2}}{r_{a}^{2}}\!(\mathbf{r}_{a}\cdot\mathbf{p}_{e})\right\rangle\bigg]-\frac{z_{e}z_{a}}{4}\langle r_{a}\bm{\varepsilon}_{a}\bm{\varepsilon}_{b}\rangle$
$\quad\pm z_{e}z_{a}\bigg\langle\imath(\varepsilon_{a}+\varepsilon_{b})^{i}\mathcal{X}^{ij}_{a}\imath\bm{\varepsilon}_{12}^{j}\bigg\rangle\bigg\}\,$, (143)
while its singular part has the form of
$S_{6}=\sum_{a}\frac{1}{8m_{e}^{2}m_{a}}\left[\langle\bm{\varepsilon}_{a}^{2}\rangle-2m_{e}\langle V_{a}^{3}\rangle+\langle V_{a}\rho_{a}\rangle\right]\,.$ (144)
C.7 evaluation of $H_{B}^{\text{1st}}$
This term generated from the transferred second-order correction $E_{B}$ ( see Eq. (30))
$H_{B}^{\text{1st}}=\{H_{B},U\}-2\langle{H_{B}}\rangle U-U(E_{0}-H_{0})U\,.$ (145)
In the evaluation, its first and third terms contain singularities.
The first term of $H_{B}^{\text{1st}}$ is expanded by inserting $U=\sum_{a}\lambda_{a}V_{a}$ (28) and $H_{B}$ (6)
$\{H_{B},U\}=\sum_{a}-\frac{\lambda_{a}}{8m_{e}^{3}}\bigg[\{\mathbf{p}_{e}^{4},V_{a}\}+m_{e}\{V_{a},\rho_{a}+\rho_{b}\}\bigg]\,.$ (146)
The matrix elements $\langle\{\mathbf{p}_{e}^{4},V_{a}\}\rangle^{(m)}$ and $\langle V_{a}\rho_{a}\rangle^{(m)}$ can be taken from Eqs. (B.2) and (86), respectively.
The third term $\langle U(E_{0}-H_{0})U\rangle$ in (145) has been considered in Sec. B.3.
Therefore, the regularized effective Hamiltonian of $H_{B}^{\text{1st}}$ is
[…]
while the corresponding singular term may be written as
$S_{B}=\sum_{a}\bigg[-\frac{1}{32m_{e}^{3}}\left(1-\frac{5m_{e}}{m_{a}}\right)\langle\bm{\varepsilon}_{a}^{2}\rangle+\frac{1}{8m_{e}^{2}}\left(1-\frac{6m_{e}}{m_{a}}\right)\langle V_{a}^{3}\rangle\bigg]\,.$ (150)
C.8 evaluation of $H_{R}^{\text{1st}}$
This term arises from the contribution of the second-order correction $E_{R}$ (see Eq. (46)):
$H_{R}^{\text{1st}}=\{H_{R},U\}-2\langle H_{R}\rangle U\,.$ (151)
Taking into account the expression of $H_{R}$, Eq. (7), the first term of $H_{R}^{\text{1st}}$ is expressed as
$\langle\{H_{R},U\}\rangle=\sum_{a}\sum_{b}-\lambda_{b}\frac{z_{e}z_{a}}{m_{e}m_{a}}\langle\{p_{e}^{i}\mathcal{W}_{a}^{ij}p_{a}^{j},V_{b}\}\rangle\,.$ (152)
In the above formula, one can see that the cross terms with $a\neq b$ are regular, while the terms with $a=b$ are singular. The regular matrix element can be evaluated as
$z_{e}z_{a}\langle\{p_{e}^{i}\mathcal{W}_{a}^{ij}p_{a}^{j},V_{b}\}\rangle_{b\neq a}$ $=2z_{e}z_{a}\left\langle p_{e}^{i}\mathcal{W}_{a}^{ij}V_{b}p_{a}^{j}\rangle-\langle\imath\bm{\varepsilon}^{i}_{b}\mathcal{W}_{a}^{ij}p_{a}^{j}\right\rangle\,.$ (153)
where the second term on the right-hand side is negligible. The divergent matrix element $\langle\{V_{a},p_{e}^{i}\mathcal{W}_{a}^{ij}p_{a}^{j}\}\rangle$ can be evaluated as
$z_{e}z_{a}\langle\{p_{e}^{i}\mathcal{W}_{a}^{ij}p_{a}^{j},V_{a}\}\rangle_{b=a}\!=\!2z_{e}z_{a}\langle p_{e}^{i}\mathcal{W}_{a}^{ij}V_{a}p_{a}^{j}\rangle\!-\!\langle\imath\varepsilon_{a}^{i}\mathcal{W}_{a}^{ij}p_{a}^{j}\rangle$
$\qquad\!-\!\langle p_{e}^{i}\mathcal{W}_{a}^{ij}\imath\varepsilon_{a}^{j}\rangle$
$=2z_{e}z_{a}\langle p_{e}^{i}\mathcal{W}_{a}^{ij}V_{a}p_{a}^{j}\rangle+\langle V_{a}\imath\bm{\varepsilon}_{a}\mathbf{p}_{e}\rangle+\langle V_{a}\imath\bm{\varepsilon}_{a}\mathbf{p}_{b}\rangle-\langle\mathbf{p}_{e}\imath\bm{\varepsilon}_{a}V_{a}\rangle\,$, (154)
where the identities of Eqs. (129) and (130) are adopted. Again, matrix element $\langle V_{a}\imath\bm{\varepsilon}_{a}\mathbf{p}_{b}\rangle$ is small enough to ignore.
Therefore, the expectation value for the regularized effective Hamiltonian is obtained by summing up all results and ignoring higher order recoil contributions
$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle=\sum_{b\neq a}\frac{z_{e}z_{a}}{2m_{e}^{2}m_{a}}\langle p_{e}^{i}\mathcal{W}_{a}^{ij}Vp_{a}^{j}\rangle-2\langle H_{R}\rangle\langle U\rangle+O(1/m_{p,d}^{2})\,$, (155)
while the singularity of $E_{R}$ can be isolated as
$S_{R}$ $=\sum_{a}-\frac{\lambda_{a}}{m_{e}m_{a}}\left[\langle V_{a}\imath\bm{\varepsilon}_{a}\mathbf{p}_{e}\rangle-\langle\mathbf{p}_{e}\imath\bm{\varepsilon}_{a}V_{a}\rangle\right]$
$\approx\sum_{a}\frac{1}{4m_{e}^{2}m_{a}}\left[\langle\bm{\varepsilon}_{a}^{2}\rangle+\langle V_{a}\rho_{a}\rangle\right]\,.$ (156)
Appendix D $m\alpha^{6}(m/M)$-order recoil corrections in hydrogen atom
Using the derivations of Sec. C  for the case of the hydrogen atom, the effective Hamiltonian for the relativistic recoil corrections of order $m\alpha^{6}(m/M)$ is obtained in the form
$H^{(6m)}_{\text{rec}}(\text{H})$ $=\frac{\bm{\varepsilon}^{2}}{8M}-\frac{V^{3}}{4M}-\frac{V\rho}{8M}+\frac{(z_{e}z_{N})^{2}}{4M}\mathbf{p}\frac{1}{r^{2}}\mathbf{p}-2\langle H_{R}\rangle U$
$\quad+\frac{(z_{e}z_{N})^{2}}{4M}\left(\mathbf{p}\frac{1}{r^{2}}\mathbf{p}+\frac{(\mathbf{p}\mathbf{r})(\mathbf{r}\mathbf{p})}{r^{4}}\right)-E_{0}H_{R}+C_{r}\frac{z_{N}^{3}}{M}\pi\delta(\mathbf{r})\,.$ (157)
where the operators $\bm{\varepsilon}^{2}$, $V^{3}$, and $V\rho$ are considered to be finite and obtained by means of the coordinate cutoff regularization. The matrix element $\langle V\rho\rangle$ is nonzero and proportional to $\langle\delta(\mathbf{r})\rangle$. The coefficient $C_{r}$ should be defined from comparison of $\langle H_{\text{rec}}^{(6m)}\rangle$ from Eq. (D) with the known result for the hydrogen $nS$ states.
The mean values of matrix elements for the nS states of hydrogen case we take from Eq. (B1) of Ref. [25] and Eqs. (23-24) of Ref. [26]. Thus the expectation value of $H_{\text{rec}}^{(6m)}$ can be expressed as
$E_{\text{rec}}^{(6m)}=\frac{m^{2}z_{N}^{6}}{Mn^{3}}\left(-\frac{23}{6}-\frac{3}{n}+\frac{35}{6n^{2}}-\frac{2}{n^{3}}\right)\,.$ (158)
This result should be compared with the analytical formula for the total $m\alpha^{6}$-order correction for the two-body system [29]. This allows to find the unkown coefficient $C_{r}$:
$C_{r}=4\ln 2-2\,$, (159)
and the last contribution in Eq. (D) includes the high-energy part (16) as well.

### 16 Mar 2026 v2

@@
 […]
+$\langle\delta H^{\prime}_{1}\rangle$ $=\frac{1}{16m_{e}^{4}}\bigg\{2E_{0}\langle p_{e}^{4}\rangle-\langle\{p_{e}^{4},V_{12}\}\rangle-\sum_{a}\bigg[\langle\{p_{e}^{4},V_{a}\}\rangle^{\prime}$
+$+\frac{4m_{e}^{2}}{m_{a}}\left(E_{0}^{2}\langle p_{a}^{2}\rangle-2E_{0}\langle Vp_{a}^{2}\rangle+\langle Vp_{a}^{2}V\rangle^{\prime}\right)\bigg]\bigg\}\,$, (38)
+$\langle\delta H^{\prime}_{2}\rangle$ $=\sum_{a}\bigg[-\frac{1}{32m_{e}^{3}}\bigg(\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{b}\rangle\mp\frac{m_{e}}{m_{a}}\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{12}\rangle\bigg)$
+$+\frac{3}{64m_{e}^{4}}\langle\{p_{e}^{2},\rho_{a}\}\rangle^{\prime}\bigg]\,$, (39)
+$\langle\delta H^{\prime}_{3}\rangle$ $=\sum_{a}\frac{1}{8m_{e}^{2}m_{a}}(\mp)\langle\bm{\varepsilon}_{a}\bm{\varepsilon}_{12}\rangle\,$, (40)
 and
@@
 [25] and Appendix A.
-$S_{\text{rec}}$ will be regularized as $R_{\text{rec}}$, namely regular part of $S_{\text{sec}}$, by eliminating singularity of high-energy contribution $H_{H}^{(6)}$.
+$S_{\text{rec}}$ will be regularized as ${(S_{\text{rec}})}_{r}$, namely regular part of $S_{\text{sec}}$,
+${(S_{\text{rec}})}_{r}=\sum_{a}\frac{1}{m_{e}^{2}m_{a}}\left(\frac{\langle\bm{\varepsilon}_{a}^{2}\rangle_{r}}{8}-\frac{m_{e}\langle V_{a}^{3}\rangle_{r}}{4}-\frac{\langle V_{a}\rho_{a}\rangle_{r}}{8}\right)\,.$ (53)
+by eliminating singularity of high-energy contribution $H_{H}^{(6)}$.
 Subsequently, the regularization of recoil contribution will rise a term determined by the $\delta$-function distribution.
 In the cut-off regularization, it is expressed as
-$H_{\text{reg}}=\sum_{a}C_{r}\frac{z_{a}^{3}}{m_{a}}\pi\delta(\mathbf{r}_{a})\,.$ (50)
+$H_{r}=\sum_{a}C_{r}\frac{z_{a}^{3}}{m_{a}}\pi\delta(\mathbf{r}_{a})\,.$ (54)
 The coefficient $C_{r}=4\ln 2-2$ is determined by comparing expectation value of regularized effective Hamiltonian with the anayltical total $m\alpha^{6}$-oder correction in hydrogen atom, see Appendix D.
 After regularization of all the contributions at $m\alpha^{6}(m/M)$ order, we obtain the effective Hamiltonians for the recoil relativistic corrections as follows:
-${E^{\prime}}^{(6)}_{\text{rel-rec}}=\langle\delta H^{\prime}_{4}+\delta H^{\prime}_{5}+\delta H^{\prime}_{6}+R_{\text{rec}}+H_{\text{reg}}+{H^{\prime}}^{\text{1st}}_{R}\rangle\,$, (51)
+${E^{\prime}}^{(6)}_{\text{rel-rec}}$ $=\langle\delta H^{\prime}_{4}+\delta H^{\prime}_{5}+\delta H^{\prime}_{6}+H_{r}+{H^{\prime}}^{\text{1st}}_{R}\rangle$
+$\quad+{(S_{\text{rec}})}_{r}\,$, (55)
 where
@@
 and
-$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle$ $=\sum_{a,b\neq a}\frac{z_{e}z_{a}}{2m_{e}^{2}m_{a}}\langle p_{e}^{i}\mathcal{W}_{a}^{ij}Vp_{a}^{j}\rangle-2\langle H_{R}\rangle\langle U\rangle\,.$ (55)
-It should be noted that the regular part $R_{\text{rec}}$ of the divergent matrix elements appeared in $S_{\text{rec}}$ (III.2) should be retained after the regularization.
+$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle$ $=\sum_{a,b\neq a}-\frac{2z_{e}z_{a}}{m_{e}^{2}m_{a}}\langle p_{e}^{i}\mathcal{W}_{a}^{ij}Up_{a}^{j}\rangle-2\langle H_{R}\rangle\langle U\rangle\,.$ (59)
+It should be noted that the regular part ${(S_{\text{rec}})}_{r}$ of the divergent matrix elements appeared in $S_{\text{rec}}$ (III.2) should be retained after the regularization.
 III.3 the regularized effective Hamiltonian
-We now sum all contributions from Eqs. (16), (19), (20), (24), (32), (50) and (51) to obtain the total corrections at orders $m\alpha^{6}$ and $m\alpha^{6}(m/M)$:
+We now sum all contributions from Eqs. (16), (19), (20), (24), (32), (54) and (III.2) to obtain the total corrections at orders $m\alpha^{6}$ and $m\alpha^{6}(m/M)$:
 $E^{(6)}$ $={E^{\prime}}_{\text{rel}}^{(6)}+{E^{\prime}}_{\text{sec}}^{(6)}+E_{\text{rad}}^{(6)}+{E^{\prime}}_{\text{rel-rec}}^{(6)}+E_{\text{rad-rec}}^{(6)}\,.$ (60)
 Comparing with previous $E^{(6)}$ of Eq.
-(25), the relativistic correction due the effective Hamiltonian $H_{\text{rel}}^{(6)}$ has been regularized and split into nonrecoil correction $E_{\text{rel}}^{(6)}$ (32) and recoil correction $E_{\text{rel-rec}}^{(6)}$ (51).
-It should be noted that the high-energy relativistic correction, so called pure recoil correction $H_{H}^{(6)}$, has been embraced in the contribution arising in the regularization $H_{\text{reg}}$ (50).
+(25), the relativistic correction due the effective Hamiltonian $H_{\text{rel}}^{(6)}$ has been regularized and split into nonrecoil correction $E_{\text{rel}}^{(6)}$ (32) and recoil correction $E_{\text{rel-rec}}^{(6)}$ (III.2).
+It should be noted that the high-energy relativistic correction, so called pure recoil correction $H_{H}^{(6)}$, has been embraced in the contribution arising in the regularization $H_{r}$ (54).
 On the other hand, the second-order corrections ${E}_{\text{sec}}^{(6)}$ is regularized as ${E^{\prime}}_{\text{sec}}^{(6)}=E^{\prime}_{B}+E_{S}+E^{\prime}_{R}$ where $E^{\prime}_{B}=\langle H^{\prime}_{B}Q(E_{0}-H_{0})^{-1}QH^{\prime}_{B}\rangle$, $E^{\prime}_{R}=2\langle H^{\prime}_{B}Q(E_{0}-H_{0})^{-1}QH_{R}\rangle$ and $E_{S}$ (24) are regular enough for numerical study.
@@
 $\langle z_{e}z_{2}p_{e}^{i}\mathcal{W}_{2}^{ij}Vp_{2}^{j}\rangle$ –1.2241 3289 5621 –1.1936 4318 5135 –1.2229 7915 8165(5) –1.1925 6219 346(2)
+$\langle z_{e}z_{1}p_{e}^{i}\mathcal{W}_{1}^{ij}Up_{1}^{j}\rangle$ 0.3795 4851 4771 0.3700 4007 8771 0.3791 6464 1365 0.3696 7653 009(1)
+$\langle z_{e}z_{2}p_{e}^{i}\mathcal{W}_{2}^{ij}Up_{2}^{j}\rangle$ 0.3784 3817 9175 0.3682 5701 6903 0.3780 2354 8783(1) 0.3678 6502 5691(6)
 $\langle z_{e}z_{1}r_{1}\bm{\varepsilon}_{1}\cdot\bm{\varepsilon}_{2}\rangle$ 0.0169 4925 9569 0.0148 8857 1832 0.0169 8530 2788 0.01492 6808 1686
@@
 Table 3: Expectation values of modified effective Hamiltonians for HD+ rovibrational states (in atomic units). The numerical values given are exact to the last digit unless otherwise noted.
-(0,0) (0,1) (1,0) (1,1)
-$\langle\delta H^{\prime}_{1}\rangle$ –1.8703 480(6) –1.8137 755(5) –1.8687 273(6) –1.8122 633(5)
-$\langle\delta H^{\prime}_{2}\rangle$ 1.7125 566(5) 1.6630 255(5) 1.7109 929(5) 1.6615 589(5)
-$\langle\delta H^{\prime}_{3}\rangle$ –0.0000 0628 5744(2) –0.0000 0603 2892(2) –0.0000 0626 5720(2) –0.0000 0601 3395(2)
-$\langle\delta{H^{\prime}}_{B}^{\text{1st}}\rangle$ 0.4020 166(1) 0.3880 342(1) 0.4016 656(1) 0.3877 119(1)
-${E^{\prime}}_{\text{rel}}^{(6)}$ 0.2442 189(8) 0.2372 782(7) 0.2439 249(8) 0.2370 001(7)
-$\langle\delta H^{\prime}_{4}\rangle$ 0.0007 150(4) 0.0007 015(4) 0.0007 144(4) 0.0007 010(4)
-$\langle\delta H^{\prime}_{5}\rangle$ 0.0005 5890 9980 0.0005 4532 3594 0.0005 5831 3896 0.0005 4475 9030
-$\langle\delta H^{\prime}_{6}\rangle$ –0.0002 1032 7438 –0.0002 0656 2370 –0.0002 1021 536(2) –0.0002 0645 96(2)
-$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle$ –0.0002 1544 0395 –0.0002 1617 5234(2) –0.0002 1533 761(1) –0.0002 1606 7579(5)
-$\langle R_{\text{rec}}\rangle$ 0.0000 3966 5983 0.0000 3761 7674 0.0000 4008 5121 0.0000 3757 6478
-$\langle H_{\text{reg}}\rangle$ 0.0004 1079 5750 0.0004 0137 2924 0.0004 1042 9068 0.0004 0102 6584
-${E^{\prime}}_{\text{rel-rec}}^{(6)}$ 0.0012 986(4) 0.0012 630(4) 0.0017 977(4) 0.0012 618(4)
+Eqs. (0,0) (0,1) (1,0) (1,1)
+$\langle\delta H^{\prime}_{1}\rangle$ (III.1) –1.8703 480(6) –1.8137 755(5) –1.8687 273(6) –1.8122 633(5)
+$\langle\delta H^{\prime}_{2}\rangle$ (39) 1.7125 566(5) 1.6630 255(5) 1.7109 929(5) 1.6615 589(5)
+$\langle\delta H^{\prime}_{3}\rangle$ (40) –0.0000 0628 5744(2) –0.0000 0603 2892(2) –0.0000 0626 5720(2) –0.0000 0601 3395(2)
+$\langle\delta{H^{\prime}}_{B}^{\text{1st}}\rangle$ (43) 0.4020 166(1) 0.3880 342(1) 0.4016 656(1) 0.3877 119(1)
+${E^{\prime}}_{\text{rel}}^{(6)}$ (32) 0.2442 189(8) 0.2372 782(7) 0.2439 249(8) 0.2370 001(7)
+$\langle\delta H^{\prime}_{4}\rangle$ (56) 0.0007 150(4) 0.0007 015(4) 0.0007 144(4) 0.0007 010(4)
+$\langle\delta H^{\prime}_{5}\rangle$ (57) 0.0005 5890 9980 0.0005 4532 3594 0.0005 5831 3896 0.0005 4475 9030
+$\langle\delta H^{\prime}_{6}\rangle$ (58) –0.0002 1032 7438 –0.0002 0656 2370 –0.0002 1021 536(2) –0.0002 0645 96(2)
+$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle$ (59) –0.0002 1544 0395 –0.0002 1617 5234(2) –0.0002 1533 761(1) –0.0002 1606 7579(5)
+${(S_{\text{rec}})}_{r}$ (53) 0.0000 3966 5983 0.0000 3761 7674 0.0000 4008 5121 0.0000 3757 6478
+$\langle H_{r}\rangle$ (54) 0.0004 1079 5750 0.0004 0137 2924 0.0004 1042 9068 0.0004 0102 6584
+${E^{\prime}}_{\text{rel-rec}}^{(6)}$ (III.2) 0.0012 986(4) 0.0012 630(4) 0.0017 977(4) 0.0012 618(4)
 Table 4: QED corrections of order $m\alpha^{6}$ for HD+ rovibrational transitions $(L,v)\rightarrow(L^{\prime},v^{\prime})$, in kHz.
@@
 [21] 1708.9(1)-1708.9(1)
+[33] 1770.95(10)--1770.95(10) [^1] The uncertainty keeps the value estimated in Ref. [21].
 The matrix elements for the operators that appear in the $m\alpha^{6}$-order correction $E^{(6m)}$, see Eq. (60), are calculated using the Hylleraas variational expansion of the wave function:
@@
 The resulting $m\alpha^{6}$-order correction to the $(0,0)\!\to\!(0,1)$ transition is determined to be $1710.684(35)~\text{kHz}$, the uncertainty of which is three orders of magnitude smaller than the value reported in previous theoretical studies [21, 23].
-As is seen, there is a discrepancy of approximately 1.8 kHz between the obtained value and the previous theory.
+As is seen, there is a discrepancy of approximately 1.8 kHz between the obtained values and the previous theory on transitions $(0,0)\!\to\!(0,1)$ and $(0,0)\!\to\!(1,1)$ as shown in Table 4.
 We attribute this to the limitations of the adiabatic Born-Oppenheimer approximation employed in previous works.
@@
 Therefore, the expectation value for the regularized effective Hamiltonian is obtained by summing up all results and ignoring higher order recoil contributions
-$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle=\sum_{b\neq a}\frac{z_{e}z_{a}}{2m_{e}^{2}m_{a}}\langle p_{e}^{i}\mathcal{W}_{a}^{ij}Vp_{a}^{j}\rangle-2\langle H_{R}\rangle\langle U\rangle+O(1/m_{p,d}^{2})\,$, (155)
+$\langle{H^{\prime}}_{R}^{\text{1st}}\rangle$ $=\sum_{b\neq a}-\frac{2z_{e}z_{a}}{m_{e}^{2}m_{a}}\langle p_{e}^{i}\mathcal{W}_{a}^{ij}Up_{a}^{j}\rangle-2\langle H_{R}\rangle\langle U\rangle$
+$\quad+O(1/m_{p,d}^{2})\,$, (159)
 while the singularity of $E_{R}$ can be isolated as